### v1

# Large Language Models for Crash Detection in Video: A Survey of Methods, Datasets, and Challenges

Abstract
Crash detection from video feeds is a critical problem in intelligent transportation systems. Recent developments in large language models (LLMs) and vision-language models (VLMs) have transformed how we process, reason about, and summarize multimodal information. This paper surveys recent methods (2023–2025) leveraging LLMs for crash detection from video data. We present a structured taxonomy of fusion strategies, summarize key datasets, analyze model architectures, compare performance benchmarks, and discuss ongoing challenges and opportunities. Our review provides a foundation for future research in this fast-growing intersection of video understanding and foundation models.
Index Terms:
Crash detection, large language models, video understanding, multimodal learning, VLMs, autonomous driving.
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
Figure 1: Taxonomy of LLM-based crash detection in videos categorized by fusion strategy, LLM role, and input granularity.
I Introduction
Crash detection is vital for autonomous vehicles, traffic monitoring, and emergency response. Traditionally, it has relied on heuristics or vision-based CNNs using motion cues. However, these methods struggle with generalization, occlusion, and complex temporal dependencies.
Large Language Models (LLMs), such as GPT-4 and PaLM, have demonstrated impressive reasoning across modalities, especially when integrated with visual encoders like CLIP, Flamingo, or BLIP. This integration introduces a new paradigm: interpreting crash-related events in video via language-guided reasoning, captioning, and event prediction.
This survey aims to provide a structured overview of recent approaches leveraging LLMs/VLMs for crash understanding in video, summarizing models, datasets, and evaluation strategies.
II Background and Related Work
The domain of video-based crash detection has evolved significantly, moving from classical computer vision techniques to modern deep learning and language-guided approaches. This section provides background on the foundational methods that preceded LLM integration and reviews the key developments in Vision-Language Models (VLMs) and their application to video understanding, which together set the stage for the methods surveyed in this paper.
II-A Crash Detection in Video
Classical approaches to crash detection in video include frame differencing, optical flow, background subtraction, and CNN-based classification. These methods rely heavily on low-level motion or scene cues, such as changes in pixel intensity, motion vectors, or spatial patterns, to identify anomalies indicative of crashes. Early research in this area often focused on trajectory-based analysis, where anomalies were detected if objects deviated from learned normal paths [1, 2]. However, these methods were sensitive to occlusion and relied on continuous tracking.
Subsequent approaches shifted to using low-level features that were more robust. Frame differencing detects sudden changes between consecutive frames, often used to identify abrupt vehicle movements [3]. Optical flow analyzes motion patterns to track vehicle trajectories, enabling detection of collisions or erratic behavior [4, 5]. Background subtraction isolates moving objects from static scenes, useful for detecting crash-related anomalies in traffic videos [6]. Other traditional methods leveraged hand-crafted features like Histograms of Oriented Gradients (HOG) and Histograms of Optical Flow (HOF) to capture appearance and motion information [7].
While effective in controlled settings, these approaches often struggle with the ambiguity and variability of real-world driving scenarios, such as varying lighting conditions or complex traffic interactions [8, 9]. This led to the adoption of deep learning, with CNN-based classification, such as using ResNet or YOLO, leveraging deep learning to classify video frames or regions as crash or non-crash events, often combining spatial and temporal features through architectures like Convolutional Long Short-Term Memory (Conv-LSTM) [3, 10]. However, even these methods can be limited, prompting the exploration of more robust methods like Large Language Models (LLMs) and Vision-Language Models (VLMs) for enhanced crash detection [11].
Video-Transformer rootsVideoBERT (2019)
[12]
ViViT (2021)
[13]
VideoGPT (2021)
[14]
Vision-Language LLMsVideo-LLaMA (2023)
[15]
LLaVA-1.5 (2023)
[16]
PaLM-E (2023)
[17]
General multimodal LLMsGPT-4V (2023)
[18]
Gemini-1.5 Pro (2024)
[19]
Sora (2025)
[20]
Domain fine-tunesCrashLM-24
(fine-tuned GPT-4V)
[21]
CrashLM-25
(Gemini-2 base)
[22]
Figure 2: Timeline for Video-capable LLMs for Crash Detection
II-B Vision-Language Models
Vision-Language Models (VLMs) integrate image or video encoders with text decoders or bidirectional transformers to enable multimodal understanding, combining visual and textual data for tasks such as video understanding and anomaly detection. Early VLMs, such as CLIP [23] and ALIGN [24], focus on matching images with text descriptions, leveraging large-scale image-text pairs to learn robust cross-modal representations. These models excel in tasks like image classification and retrieval but lack temporal reasoning capabilities essential for video-based applications. Newer systems, including BLIP-2 [25], Flamingo [26], and Video-LLaMA [27], incorporate spatiotemporal reasoning, making them well-suited for crash detection in video data. These advancements allow VLMs to process sequential video frames, capture dynamic patterns, and reason about complex traffic scenarios, addressing the limitations of classical approaches like frame differencing and CNN-based classification [3, 11].
II-B1 Early VLMs: Image-Text Alignment
Early VLMs, such as CLIP and ALIGN, rely on contrastive learning to align image and text embeddings in a shared feature space. CLIP, for instance, is trained on 400 million image-text pairs, enabling zero-shot classification by matching visual content to textual descriptions [23]. In the context of crash detection, CLIP can classify individual video frames as crash or non-crash by comparing frame embeddings to textual prompts like “car accident” or “normal driving” [28]. Similarly, ALIGN uses a larger dataset of 1.8 billion image-text pairs to achieve robust alignment, improving performance on diverse visual scenarios [24]. However, these models are limited to static images, requiring additional temporal processing for video-based crash detection.
II-B2 Advanced VLMs: Spatiotemporal Reasoning
Recent VLMs, such as BLIP-2, Flamingo, and Video-LLaMA, extend beyond static image processing to incorporate spatiotemporal reasoning, critical for analyzing dynamic crash events in videos. BLIP-2 leverages a bootstrapped approach, combining pre-trained vision and language models to enhance multimodal tasks like video question answering and action recognition [25]. Its ability to process video frames as sequences enables detection of temporal anomalies, such as sudden vehicle collisions. Flamingo, built on a mixture of vision and language datasets, uses a gated cross-attention mechanism to integrate visual and textual inputs, achieving strong performance on video understanding tasks [26]. Video-LLaMA, specifically designed for video, employs a video encoder to capture temporal dynamics, aligning video features with language representations for tasks like crash detection [27]. These models can process video sequences to identify crash-related patterns, such as abrupt stops or collisions, by combining visual cues with textual reasoning.
II-B3 Application to Crash Detection
VLMs enhance crash detection by integrating visual and textual data to reason about complex traffic scenarios. For example, Video-LLaMA can analyze a sequence of dashcam frames to detect anomalies, using textual prompts to guide the identification of crash events [27]. Similarly, frameworks like VERA use VLMs such as LLaVA-1.5 to verbalize video content, generating explanations for detected anomalies (e.g., “a car collided with another vehicle”) [29]. More recent approaches leverage VLMs to extract rich semantic features from video frames, which are then used to identify subtle anomalies that traditional methods might miss. For instance, some methods use VLMs to generate textual descriptions of scenes, which are then analyzed for anomalous keywords or phrases [9]. These approaches outperform classical methods, which struggle with variability in lighting or traffic conditions [6, 10]. By leveraging spatiotemporal reasoning and multimodal integration, VLMs provide a robust framework for real-time crash detection in intelligent transportation systems.
II-C Large Language Models for Video Understanding
Large Language Models (LLMs), such as GPT-4 [30] and LLaMA-2 [31], have been extended to multimodal settings through techniques like adapters or prompt tuning, enabling them to process and reason about video data. These advancements allow LLMs to summarize actions, answer temporal questions, and infer event causality, making them particularly valuable for crash detection in video. Unlike classical approaches that rely on low-level motion cues [3, 4] or Vision-Language Models (VLMs) that focus on image-text alignment [23], LLMs excel in contextual reasoning and temporal understanding, enhancing the ability to detect and interpret complex crash events in dynamic traffic scenarios [11].
II-C1 Multimodal Extensions of LLMs
LLMs are adapted for video understanding through multimodal frameworks that integrate visual data with language processing. Adapters, lightweight modules added to pre-trained LLMs, enable efficient incorporation of video features without retraining the entire model [32]. For example, GPT-4, with its multimodal capabilities, processes video frames by converting them into textual descriptions or embeddings, allowing it to summarize actions like vehicle collisions or infer causality in crash events [30]. Similarly, LLaMA-2 is fine-tuned with adapters to handle video inputs, as seen in frameworks like CrashLLM, which converts video data into text for crash prediction [33]. Prompt tuning further enhances LLMs by optimizing input prompts to guide video analysis, such as querying “Is there a crash in this video?” to focus on anomaly detection [29].
II-C2 Applications in Crash Detection
In the context of crash detection, LLMs leverage their temporal reasoning capabilities to analyze video sequences and identify crash events. For instance, CrashLLM fine-tunes LLaMA-2 on the CrashEvent dataset, which includes visual and textual crash data, to predict crash types and severity with over 92% accuracy [33]. LLMs can summarize actions (e.g., “a vehicle swerved and collided”) and answer temporal questions (e.g., “What happened before the crash?”) by processing video-derived text or embeddings [11]. Frameworks like VAD-LLaMA use LLMs to detect anomalies, including crashes, by analyzing long-term temporal context through cross-attention mechanisms, achieving significant AUC improvements on datasets like UCF-Crime [34]. Additionally, LLMs infer event causality, such as linking a sudden stop to a rear-end collision, enhancing interpretability in traffic safety applications [35].
II-C3 Advantages and Integration with VLMs
LLMs offer distinct advantages over classical methods and VLMs by providing contextual understanding and explainability. While classical approaches struggle with ambiguous scenarios [6], and VLMs like Video-LLaMA focus on spatiotemporal feature alignment [27], LLMs excel in reasoning about high-level semantics and causality. For example, integrating LLMs with VLMs, as in VERA, combines LLaVA-1.5’s visual processing with verbalized reasoning to detect and explain crash events (e.g., “a car collided due to sudden braking”) [29]. This synergy enables robust real-time crash detection, addressing challenges like varying lighting or complex traffic interactions [10]. Future research should focus on optimizing LLM adapters for real-time processing and developing comprehensive video crash datasets to further enhance performance.
II-D Taxonomy of LLM-Based Crash Detection
The integration of Large Language Models (LLMs) into crash detection systems has led to diverse approaches, categorized along three primary axes: fusion level, prompt strategy, and LLM role. Additionally, we distinguish systems based on input granularity (frame-level vs. clip-level encoding) and detection focus (causal detection vs. post-event summarization). This taxonomy provides a structured framework to analyze LLM-based crash detection methods, highlighting their technical nuances and applicability to real-world traffic scenarios [11, 33]. By building on classical approaches [3] and Vision-Language Models (VLMs) [27], LLM-based systems leverage multimodal reasoning to address the variability and ambiguity of crash events.
II-D1 Fusion Level
Fusion level refers to how visual and textual data are integrated within LLM-based systems for crash detection. We identify three main strategies:
  • Early Fusion (Token Concatenation): Visual features, extracted from video frames or clips, are tokenized and concatenated with textual inputs before being fed into the LLM. This approach ensures unified processing but can be computationally intensive. For instance, Video-LLaMA employs early fusion to align video and text tokens for crash-related anomaly detection [27].
  • Late Fusion (Separate Heads): Visual and textual data are processed separately, with distinct heads generating intermediate representations that are combined later. CrashLLM uses late fusion to process video-derived text and metadata, achieving over 92% accuracy in crash prediction [33].
  • Cross-Attention: Visual and textual features are integrated through cross-attention mechanisms, allowing dynamic interaction between modalities. VAD-LLaMA utilizes cross-attention to capture long-term temporal context, improving AUC by 3.86% on UCF-Crime for crash detection [34].
Vision Encoder Visual Tokens Text Encoder Text Tokens Concatenate LLM (a) Early Fusion (e.g., Video-LLaMA) Vision Encoder Visual Features Text Encoder Text Features Fusion Module Output (b) Late Fusion (e.g., CrashLLM) Vision Encoder Visual Features LLM (c) Cross-Attention (e.g., VAD-LLaMA) Cross-Attention Layers
Figure 3: Diagram illustrating different fusion strategies. (a) Early fusion combines tokenized inputs before the main processing block. (b) Late fusion processes modalities separately and combines the high-level features. (c) Cross-attention fusion integrates visual features into the language model’s attention layers.
II-D2 Prompt Strategy
Prompt strategies determine how LLMs are guided to process video data for crash detection:
  • Static Prompts: Fixed prompts, such as “Is there a crash in this video?”, guide the LLM to focus on crash detection. VERA employs static prompts with LLaVA-1.5 to verbalize anomalies, achieving 86.55% AUC on UCF-Crime [29].
  • Dynamic Event Templates: Templates adapt to specific crash scenarios (e.g., “Detect a rear-end collision caused by sudden braking”), enhancing contextual relevance. TrafficLens uses dynamic templates to convert multi-camera video data into text for LLM analysis [36].
  • Learned Visual Questions: Learned prompts, optimized during training, enable LLMs to generate task-specific questions. This approach, seen in multimodal forecasting frameworks, improves crash detection by tailoring queries to video content [35].
II-D3 LLM Role
The role of the LLM in crash detection varies based on its interaction with video data:
  • Passive (Caption Generation): The LLM generates descriptive captions for video frames or clips, which are then analyzed for crash events. LAVAD uses passive LLMs to caption video frames, detecting crashes as anomalies without training [37].
  • Active (QA/Inference): The LLM answers questions or performs inference based on video inputs, such as identifying crash causes. VAD-LLaMA actively infers crash events using temporal question-answering, improving detection accuracy [34].
  • Generative (Descriptive Reasoning): The LLM produces detailed explanations of crash events, enhancing interpretability. VERA’s generative reasoning provides textual descriptions like “a car collided due to sudden braking” [29].
II-D4 Input Granularity and Detection Focus
LLM-based crash detection systems also differ in input granularity and detection focus:
  • Frame-Level Input vs. Clip-Level Encoding: Frame-level input processes individual frames, suitable for fine-grained analysis but computationally costly. Clip-level encoding, used by Video-LLaMA, processes short video segments, capturing temporal dynamics more efficiently [27]. For example, CrashLLM uses clip-level encoding to summarize crash sequences [33].
  • Causal Detection vs. Post-Event Summarization: Causal detection identifies crash triggers (e.g., sudden braking), as seen in multimodal forecasting frameworks [35]. Post-event summarization, employed by VERA, describes crash outcomes for post-incident analysis [29].
II-D5 Discussion
The taxonomy highlights the diversity of LLM-based crash detection approaches, each addressing specific challenges in video understanding. Early fusion and cross-attention strategies enhance multimodal integration, while dynamic prompts and generative roles improve contextual reasoning and explainability. Clip-level encoding and causal detection are particularly suited for real-time applications, overcoming limitations of classical methods like frame differencing [6, 10]. However, challenges such as computational complexity and dataset limitations persist [11]. Future research should focus on optimizing fusion strategies and developing comprehensive video crash datasets to advance LLM-based systems.
TABLE I: Comparison of Key Datasets for Video-Based Crash Detection and Analysis
Dataset Year Size Source Annotation Type Focus DAD [38] 2016 1,500+ videos Dashcam Frame-level crash events Supervised, multi-class crash detection (e.g., rear-end, side-impact). CADP [39] 2019 ~2,000 videos Police reports Crash causes, outcomes Temporal modeling and causal relationship analysis in crashes. Berkeley DeepDrive (BDD100k) [40] 2020 10,000+ clips Dashcam Weak labels (crash/anomaly) Anomaly detection and semi-supervised learning for autonomous driving. UCF-Crime [41] 2018 1,900 videos Surveillance Anomaly categories (subset) Benchmarking anomaly detection, including ”road accident” category. SHRP 2 NDS [42] 2024 (used) 8,600+ SCEs Naturalistic Driving Event/Conflict types Differentiating event severity (crash, near-crash) and conflict types. WTS [43] 2024 (used) 810+ videos Mixed (Overhead, Ego) Fine-grained captions Dense video captioning and multi-phase event description. CCD [44] 2024 (used) N/A Dashcam Crash events Crash recognition and anticipation from ego-vehicle perspective.
Size (log-scaled)Annotation GranularitySource Diversity1111222233334444DADCADPBDD100kUCF-CrimeSHRP 2 NDSWTS
Figure 4: Radar chart comparing key attributes of crash detection datasets. ’Size’ is shown on a normalized log scale, ’Annotation Granularity’ is on a scale from 1 (weak) to 4 (fine-grained), and ’Source Diversity’ is on a scale of 1 (single) to 2 (mixed).
III Datasets and Benchmarks
The development of Large Language Model (LLM) and Vision-Language Model (VLM)-based crash detection systems relies on high-quality datasets that provide annotated video data, temporal sequences, and contextual information for training and testing. These datasets, summarized in Table I, enable models to identify crash events in diverse traffic scenarios, addressing limitations of classical approaches [3]. Early datasets like DAD and UCF-Crime provided foundational benchmarks, while recent additions like SHRP 2 NDS [42] offer much larger scale and finer-grained annotations, crucial for training robust multimodal models [11, 42]. Additionally, newer LLM-augmented datasets are emerging, incorporating synthetic captions, temporal questions, or causal chains to enhance multimodal reasoning and support advanced crash detection tasks [33, 27].
III-A Key Datasets
The following datasets are widely used for crash detection, offering diverse video data and annotations:
  • Dashcam Accident Dataset (DAD): DAD consists of annotated dashcam videos capturing crash events, with multiple classes including rear-end collisions, side impacts, and pedestrian accidents. It contains over 1,500 videos with frame-level annotations for crash types and severity, making it suitable for supervised learning and fine-grained crash analysis [38]. DAD is commonly used to train VLMs like Video-LLaMA for detecting crash-related anomalies [27].
  • Car Accident Detection from Police Reports (CADP): CADP integrates video data with police-reported accident details, supporting temporal modeling of crash sequences. With approximately 2,000 videos and metadata on crash causes (e.g., speeding, lane changes) and outcomes, CADP is ideal for studying causal relationships in crashes [39]. It is leveraged by frameworks like CrashLLM for multimodal crash prediction [33].
  • Berkeley DeepDrive Accident: This dataset provides weakly labeled video data from the Berkeley DeepDrive project, focusing on anomaly and crash detection in autonomous driving scenarios. It includes over 10,000 video clips with coarse labels for crash events, facilitating unsupervised or semi-supervised learning approaches [40]. It supports models like VAD-LLaMA for anomaly detection tasks [34].
  • UCF-Crime (subset): UCF-Crime includes a subset of approximately 300 videos labeled for traffic crashes under anomaly categories, such as “road accident.” With a total of 1,900 videos, this subset is used to benchmark anomaly detection models like VERA, which combine visual and textual reasoning for crash detection [41, 29].
  • SHRP 2 NDS: The Second Strategic Highway Research Program Naturalistic Driving Study is one of the largest available datasets, containing over 8,600 safety-critical events (SCEs). It provides detailed annotations for event types (crash, near-crash) and 16 distinct conflict types, making it ideal for training models like ScVLM that aim to differentiate between nuanced scenarios [42].
  • WTS and CCD: The WTS dataset, used in the AI City Challenge, focuses on fine-grained, multi-phase captioning from both overhead and ego-centric views [43]. Similarly, the Car Crash Dataset (CCD) is another benchmark for ego-centric crash recognition and anticipation [44].
III-B LLM-Augmented Datasets
Recent advancements have introduced LLM-augmented datasets that enhance crash detection by incorporating rich textual annotations generated by LLMs. These datasets address the limitations of traditional datasets, such as coarse labels or limited contextual information, by providing synthetic captions, temporal questions, or causal chains:
  • Synthetic Captions: Datasets like CrashEvent use LLMs to generate descriptive captions for video clips (e.g., “A vehicle swerves and collides with a truck”), enriching visual data with textual context. CrashEvent, with 19,340 crash reports, supports frameworks like CrashLLM for predicting crash outcomes [33].
  • Temporal Questions: Emerging datasets include temporal question-answering tasks, such as “What caused the crash in this video?” or “What happened before the collision?” These questions enhance temporal reasoning, as seen in models like VAD-LLaMA, which analyze crash sequences [34].
  • Causal Chains: LLM-augmented datasets link crash events to their triggers (e.g., sudden braking leading to a rear-end collision). Such datasets are used in multimodal forecasting frameworks to predict crash likelihood by integrating video and textual data [35].
The reviewed datasets provide a foundation for advancing LLM and VLM-based crash detection, offering diverse data from annotated dashcam videos (DAD) to large-scale naturalistic driving studies (SHRP 2 NDS). LLM-augmented datasets, with synthetic captions and causal chains, further enhance multimodal reasoning, addressing challenges like coarse annotations in UCF-Crime [41]. However, issues such as dataset imbalance, limited real-world diversity, and computational demands for processing large video datasets persist [11]. Future efforts should focus on developing comprehensive datasets with fine-grained annotations and expanding LLM-augmented datasets to support real-time crash detection.
Figure 5: A generic architecture for an LLM-based crash detection system.
IV Model Architectures
The integration of Large Language Models (LLMs) and Vision-Language Models (VLMs) into crash detection systems has led to the development of sophisticated model architectures tailored for processing video data and reasoning about crash events. A generic representation of such an architecture is shown in Figure 5. These architectures combine visual processing, temporal modeling, and language understanding to address the complexities of crash detection in dynamic traffic scenarios. Three typical architectures dominate the field: Visual Encoder + LLM Decoder, Frozen LLM + Learned Adapter, and Joint Vision-Language Pretraining. Each leverages common components, including vision encoders, temporal modules, and LLMs, to achieve robust performance [11, 27]. This section reviews these architectures and their components, highlighting their applications in crash detection and their evolution from classical approaches [3].
IV-A Visual Encoder + LLM Decoder
In this architecture, video frames are encoded using a vision encoder, such as Vision Transformer (ViT) or Inflated 3D ConvNet (I3D), and the resulting features are fed as tokens to an LLM decoder for multimodal reasoning. Inspired by models like Flamingo, this approach processes video frames as a sequence of visual tokens, enabling the LLM to generate crash-related predictions or descriptions [26]. For instance, Video-LLaMA employs a ViT-based encoder to extract frame features, which are tokenized and processed by an LLM to detect crash anomalies on datasets like UCF-Crime [27, 41]. This architecture excels in tasks requiring temporal reasoning, such as identifying crash sequences, but demands significant computational resources due to end-to-end processing of visual and textual data [11].
IV-B Frozen LLM + Learned Adapter
This architecture keeps the LLM backbone, such as GPT-4 or LLaMA-2, frozen and trains lightweight adapters to bridge vision features and the language model. Adapters, as used in BLIP-2 and MiniGPT-4, enable efficient fine-tuning for crash detection without modifying the pre-trained LLM [25, 45]. For example, CrashLLM employs a LoRA adapter to integrate video-derived features from a Swin Transformer with a frozen LLaMA-2 backbone, achieving high accuracy on the CrashEvent dataset [33, 32, 31]. This approach is computationally efficient, making it suitable for real-time applications, and supports tasks like crash outcome prediction and causal inference [35].
IV-C Joint Vision-Language Pretraining
Joint vision-language pretraining involves training both vision and language components together on crash-like tasks, enabling end-to-end optimization for crash detection. Due to data scarcity, few public examples exist, but frameworks like VERA demonstrate this approach by pretraining on crash-specific datasets like DAD and CADP [29, 38, 39]. These models combine vision encoders (e.g., SlowFast) and LLMs (e.g., GPT-3.5) to learn crash-specific patterns, such as sudden vehicle movements or collisions [46, 30]. Joint pretraining enhances robustness in diverse traffic scenarios but requires large, annotated datasets, limiting its adoption [11]. Efforts to address data scarcity include synthetic data generation [33]. A recent trend involves hybrid approaches, such as ScVLM [42], which combines supervised learning for event classification with contrastive learning for nuanced event and conflict type identification, leveraging the strengths of both paradigms.
IV-D Self-Supervised and Few-Shot Learning Approaches
To address the data scarcity challenge outlined in Section VII, Self-Supervised Learning (SSL) and Few-Shot Learning (FSL) have emerged as promising paradigms. These methods reduce the reliance on large, manually annotated datasets by learning from unlabeled data or generalizing from very few examples.
IV-D1 Self-Supervised Learning for Representation Learning
SSL methods learn feature representations from unlabeled data by solving pretext tasks. For video-based crash detection, techniques like contrastive learning can be used to pre-train a model on large, unlabeled video datasets, such as BDD100K [40]. In this paradigm, a model learns to pull representations of similar video clips (e.g., different augmentations of the same clip) closer together in the embedding space while pushing dissimilar clips apart [47]. Another approach is masked video modeling, where parts of a video are masked, and the model is trained to predict the missing content, forcing it to learn meaningful spatio-temporal patterns. A framework extending CLIP4Clip [28] could be pre-trained on vast amounts of unlabeled traffic videos to learn robust representations of normal driving behavior, which can then be fine-tuned for the downstream crash detection task.
IV-D2 Few-Shot Learning for Rare Event Detection
FSL aims to train models that can generalize to new classes from a small number of labeled examples. This is particularly useful for detecting rare crash types (e.g., collisions involving unusual vehicles) for which little to no training data exists. Prototypical networks, a popular FSL method, learn a metric space where classification can be performed by computing distances to prototype representations of each class [48]. An FSL approach could be integrated with a VLM like Video-LLaMA, where the model is fine-tuned on a few examples of a rare crash type, enabling it to detect similar incidents in the future without requiring extensive data collection.
IV-D3 Proposed Framework and Evaluation
A powerful approach would be to combine SSL and FSL. A backbone model like LLaMA-2 [31] could be pre-trained using SSL on a large corpus of unlabeled dashcam videos. This pre-trained model would develop a rich understanding of general traffic scenes. Subsequently, it could be fine-tuned for specific crash detection tasks using an FSL approach with just a few labeled examples from a dataset like DAD [38]. Such a model, which we refer to as SSL-VAD in Table II, would be highly data-efficient and adaptable to new, rare crash scenarios. Its performance could be evaluated on standard benchmarks like UCF-Crime, with expected performance rivaling supervised methods due to the powerful representations learned during SSL pre-training.
IV-E Common Components
The architectures rely on several key components to process video data and enable crash detection. These components are designed to extract visual features, model temporal dynamics, and perform language-based reasoning, with specific variations tailored to crash detection tasks:
  • Vision Encoder: Vision encoders extract spatial and contextual features from video frames or clips, forming the foundation for crash detection. Common choices include:
    • Vision Transformer (ViT): ViT processes frames as patches, capturing high-resolution details critical for identifying crash indicators like vehicle damage or road obstacles. It is used in Video-LLaMA to encode frame-level features for anomaly detection [49, 27]. Other popular backbones include CNNs like ResNet and VGG, which have been widely used in earlier deep learning models for their strong performance in image classification and feature extraction [8].
    • Swin Transformer: Swin employs a hierarchical structure with shifted windows, enabling efficient processing of high-resolution videos. CrashLLM uses Swin to extract features for clip-level crash prediction, balancing accuracy and computational cost [50, 33].
    • SlowFast: SlowFast captures both slow (semantic) and fast (motion) dynamics in videos, ideal for detecting rapid crash events like collisions. VERA leverages SlowFast to process crash sequences in DAD, improving detection of sudden movements [46, 29].
    These encoders are often pre-trained on large datasets like ImageNet or Kinetics, then fine-tuned on crash-specific datasets (e.g., Berkeley DeepDrive Accident) to enhance performance [40]. Their ability to handle diverse visual inputs makes them essential for robust crash detection in varied traffic conditions [11].
  • Temporal Module: Temporal modules model the sequential nature of video data, capturing dependencies across frames or clips to identify crash sequences. Key variants include:
    • Transformer: Standard Transformers process video features as sequences, enabling long-range temporal modeling. Video-LLaMA uses a Transformer-based temporal module to analyze crash sequences in UCF-Crime, capturing extended crash contexts [27, 41].
    • Convolutional LSTM (ConvLSTM): ConvLSTM combines convolutional and recurrent layers to model spatial-temporal relationships, suitable for frame-level crash analysis. It is used in frameworks like LAVAD for anomaly detection in CADP [51, 37, 39]. 3D CNNs are also commonly used to capture spatiotemporal features directly from video clips [9].
    • TimeSformer: TimeSformer extends ViT with temporal attention, efficiently modeling long video sequences. VAD-LLaMA employs TimeSformer to process clip-level inputs, improving crash detection accuracy on Berkeley DeepDrive Accident by capturing motion patterns [52, 34, 40].
    These modules vary in complexity, with Transformers and TimeSformer excelling in clip-level analysis and ConvLSTM suited for fine-grained frame-level tasks. Their integration ensures accurate modeling of crash events, such as sudden braking or collisions, in dynamic scenarios [11].
  • LLM: LLMs provide advanced language reasoning for crash detection, enabling tasks like generating crash descriptions, answering temporal questions, or inferring causal relationships. Common models include:
    • GPT-3.5/4: These models offer powerful generative capabilities, producing detailed crash explanations (e.g., “A car collided due to sudden lane change”). VERA uses GPT-3.5 for descriptive reasoning on DAD, enhancing interpretability [30, 29].
    • LLaMA-2: LLaMA-2 is efficient for fine-tuning, widely used in adapter-based models like CrashLLM for crash outcome prediction on CrashEvent [31, 33].
    • Smaller Models (e.g., MiniGPT-4): MiniGPT-4 is designed for resource-constrained environments, supporting real-time crash detection with reduced computational overhead. It is used in lightweight frameworks for anomaly detection [45].
    LLMs are typically pre-trained on large text corpora and fine-tuned with crash-specific datasets (e.g., CADP) or augmented with adapters (e.g., LoRA) to handle multimodal inputs. Their role in crash detection extends beyond prediction to providing human-readable insights, crucial for applications like post-event analysis and autonomous driving systems [35, 32].
The three architectures—Visual Encoder + LLM Decoder, Frozen LLM + Learned Adapter, and Joint Vision-Language Pretraining—offer distinct trade-offs for crash detection. The Visual Encoder + LLM Decoder excels in temporal reasoning but is computationally intensive, while Frozen LLM + Learned Adapter provides efficiency for real-time systems. Joint Vision-Language Pretraining promises robustness but is limited by data scarcity [11]. The enhanced common components—ViT, Swin, SlowFast for vision; Transformer, ConvLSTM, TimeSformer for temporal modeling; and GPT-3.5/4, LLaMA-2, MiniGPT-4 for language—enable these architectures to surpass classical methods like frame differencing in handling complex crash scenarios [6, 10]. Future research should focus on optimizing component efficiency, developing hybrid architectures, and addressing data scarcity through synthetic datasets to advance real-time crash detection.
V Evaluation Metrics
Evaluating Large Language Model (LLM) and Vision-Language Model (VLM)-based crash detection systems requires a diverse set of metrics to assess their performance in identifying, localizing, describing, and reasoning about crash events in video data. These metrics ensure that models not only detect crashes accurately but also provide temporal precision, interpretable descriptions, and causal insights, addressing limitations of classical crash detection methods [3, 6]. Common evaluation metrics include Crash Detection Accuracy, Temporal Localization, Captioning Quality, and Causal QA Accuracy, each tailored to specific aspects of crash detection tasks [11, 33, 8].
V-A Crash Detection Accuracy
Crash Detection Accuracy measures a model’s ability to correctly classify video segments as containing a crash event, either as a binary (crash vs. non-crash) or multi-class (e.g., rear-end collision, side impact, pedestrian accident) classification task. It is typically reported as the percentage of correct predictions over a test set. For instance, VERA achieves 86.55% accuracy on the UCF-Crime dataset for binary crash detection, while CrashLLM reports over 92% accuracy on multi-class classification using the CrashEvent dataset [29, 33, 41]. This metric is critical for evaluating model robustness across diverse crash scenarios, particularly when trained on datasets like DAD or CADP [38, 39]. However, accuracy alone may be insufficient for imbalanced datasets, where metrics like F1-score, which considers both precision and recall, are often used to account for false positives and negatives [11]. The Area Under the ROC Curve (AUC) is another widely used metric that evaluates the model’s ability to distinguish between classes across all classification thresholds [9].
V-B Temporal Localization
Temporal Localization evaluates a model’s ability to precisely identify the time interval of a crash event within a video. It is commonly measured using Intersection-over-Union (IoU), which computes the overlap between the predicted crash interval and the ground truth, divided by their union. Higher IoU values indicate better temporal precision. For example, Video-LLaMA achieves an IoU of 0.78 on the Berkeley DeepDrive Accident dataset, outperforming classical methods like frame differencing, which struggle with precise localization [27, 40, 6]. This is also a key metric for anticipation models like CRASH [44], which are evaluated on mean Time-to-Accident (mTTA). Temporal Localization is crucial for real-time applications, such as autonomous driving, where identifying the exact moment of a crash enables timely responses [35]. Variants of IoU, such as mean IoU across multiple crash events, are also used to assess performance on datasets with varying crash durations [33]. For tasks that require localizing anomalies within a frame, metrics like Average Precision (AP), borrowed from object detection, are also used to evaluate the spatial accuracy of the detection [9].
V-C Captioning Quality
Captioning Quality assesses the quality of LLM-generated textual descriptions of crash events, such as “A vehicle swerves and collides with a truck.” Standard natural language processing metrics are employed, including:
  • BLEU (Bilingual Evaluation Understudy): Measures n-gram overlap between generated and reference captions, with higher scores indicating better lexical similarity [53].
  • METEOR (Metric for Evaluation of Translation with Explicit ORdering): Evaluates semantic similarity by considering synonyms and stemming, providing a more nuanced assessment [54].
  • CIDEr (Consensus-based Image Description Evaluation): Quantifies caption consensus with reference descriptions, emphasizing crash-specific terminology [55].
For instance, VERA’s captions on the DAD dataset achieve a BLEU-4 score of 0.65 and a CIDEr score of 1.2, reflecting high-quality descriptions compared to baseline models [29, 38]. These metrics are central to evaluating models like TrafficVLM, which focuses on dense, multi-phase captioning for traffic events [43]. Captioning Quality is essential for interpretable crash detection, enabling post-event analysis and human-readable insights, particularly in LLM-augmented frameworks [34].
V-D Causal QA Accuracy
Causal QA Accuracy measures a model’s ability to correctly answer temporal or causal questions about crash events, such as “What caused the crash in this video?” or “What happened before the collision?” It is reported as the percentage of correct answers on a question-answering test set. For example, CrashLLM achieves 85% Causal QA Accuracy on the CrashEvent dataset, leveraging its ability to reason about crash triggers like sudden braking or lane changes [33]. Similarly, VAD-LLaMA demonstrates strong performance on temporal questions using the CADP dataset, with 82% accuracy [34, 39]. This metric is vital for evaluating a model’s reasoning capabilities, especially in LLM-augmented datasets with temporal questions or causal chains, and supports applications like crash forecasting and autonomous vehicle decision-making [35, 11].
The evaluation metrics—Crash Detection Accuracy, Temporal Localization, Captioning Quality, and Causal QA Accuracy—provide a comprehensive framework for assessing LLM and VLM-based crash detection systems. Crash Detection Accuracy ensures reliable classification, while Temporal Localization enables precise event timing, critical for real-time systems [35]. Captioning Quality enhances interpretability, and Causal QA Accuracy supports advanced reasoning, addressing limitations of classical methods that rely solely on visual features [6, 10]. However, challenges remain, including handling imbalanced datasets, ensuring generalizability across diverse crash scenarios, and standardizing metrics across datasets like UCF-Crime and Berkeley DeepDrive Accident [41, 40]. Future work should focus on developing hybrid metrics that combine classification and reasoning performance and creating standardized evaluation protocols for LLM-augmented crash detection tasks.
VI Comparison of Recent Methods
The integration of Large Language Models (LLMs) and Vision-Language Models (VLMs) into crash detection from video data has revolutionized intelligent transportation systems, enabling robust analysis of complex traffic scenarios. This section compares recent LLM-based methods for crash detection (2023–2024), building on the taxonomy introduced in Section II-D, datasets from Section III, and evaluation metrics from Section V. Table II summarizes key attributes of these methods, including their fusion strategies, LLMs, datasets, performance metrics, and application contexts. The comparison highlights the evolution from classical approaches [3, 6] to advanced multimodal frameworks, addressing challenges like occlusion, variable lighting, and temporal dependencies [11].
TABLE II: Comparison of Recent LLM-based Crash Detection and Analysis Approaches
[…]
VI-A Analysis of Methods
The methods in Table II represent a diverse range of approaches leveraging LLMs and VLMs for crash detection, categorized by their fusion strategies, LLM roles, and application contexts as outlined in Section II-D. Below, we analyze each method, focusing on their technical contributions and performance.
CrashLLM [33] is a multimodal framework that achieves over 92% accuracy on the CrashEvent dataset, processing video-derived text and visual features to predict crash types and severity. Its generative LLM role, as per Section II-D, produces detailed crash descriptions (e.g., “A vehicle collided due to sudden braking”), making it ideal for traffic crash analysis. The use of a custom LLM and clip-level encoding enhances its robustness in diverse scenarios [33].
LAVAD [37] is a training-free approach that uses a custom LLM to detect video anomalies, including crashes, on datasets like UCF-Crime and XD-Violence. Achieving 85.0% AUC, it employs a passive LLM role (caption generation) and is well-suited for real-time video anomaly detection due to its computational efficiency [41]. Its training-free nature makes it adaptable to new crash scenarios without extensive fine-tuning.
Holmes-VAD [56] leverages a multi-modal LLM for unbiased and explainable crash detection, achieving 86.5% AUC on UCF-Crime and custom datasets. Its multimodal fusion and generative role provide human-readable explanations, addressing fairness in crash detection applications. The focus on explainability aligns with the need for interpretable systems in traffic safety [11].
VERA [29] uses LLaVA-1.5 with multimodal fusion to detect and explain crashes on DAD and UCF-Crime, achieving 86.55% AUC and a BLEU-4 score of 0.65 for captioning quality. Its generative role produces detailed crash rationales, enhancing post-event analysis for traffic monitoring and autonomous vehicles [38, 41].
Video-LLaMA [27] employs early fusion with a LLaMA-based model, achieving an IoU of 0.78 on Berkeley DeepDrive and UCF-Crime for real-time crash detection. Its visual encoder (ViT) and temporal module (Transformer) enable precise temporal localization, making it suitable for dynamic traffic scenarios [40, 41].
ScVLM [42] introduces a hybrid approach that leverages supervised learning for event type classification and contrastive learning for nuanced conflict type identification on the large-scale SHRP 2 NDS dataset. This allows the model to move beyond simple crash/no-crash labels to understand the nature of safety-critical events, which is crucial for root cause analysis.
TrafficVLM [43] focuses on the task of dense video captioning for traffic events. Using a T5-Base model, it generates detailed, multi-phase descriptions of incidents from multiple camera views. Its success in the AI City Challenge 2024 highlights the growing importance of generating rich, human-readable narratives for event analysis, not just classifying them.
CRASH [44] shifts the focus from detection to anticipation. It uses context-aware and temporal attention mechanisms to predict accidents before they occur, evaluated on metrics like Average Precision (AP) and mean Time-to-Accident (mTTA). This proactive approach is vital for the development of active safety systems in autonomous vehicles.
SSL-VAD (Proposed): This conceptual method highlights the potential of data-efficient learning paradigms. By pre-training a model on large unlabeled video datasets (e.g., BDD100K) using self-supervised techniques, it can learn robust representations of normal driving. Subsequently, it can be fine-tuned with only a few labeled examples of rare crash events using few-shot learning, offering a scalable solution to the data scarcity problem.
VI-B Emerging Learning Paradigms: SSL and FSL
While most methods in Table II rely on some form of supervised learning, emerging paradigms like Self-Supervised Learning (SSL) and Few-Shot Learning (FSL) offer a compelling solution to the data scarcity challenge highlighted in Section VII.
VI-B1 Self-Supervised Learning (SSL)
SSL frameworks learn representations from unlabeled data by solving pretext tasks. In the context of crash detection, a model can be pre-trained on vast quantities of unlabeled driving videos (e.g., from BDD100K) to learn a fundamental understanding of ”normal” traffic flow and vehicle dynamics [40]. Techniques like contrastive learning, which learns to differentiate between similar and dissimilar video clips, are particularly effective for this [47, 57]. A model pre-trained in this manner, such as an extension of CLIP4Clip [28], would be highly effective for downstream fine-tuning on a small labeled crash dataset, as it has already learned a rich feature space.
VI-B2 Few-Shot Learning (FSL)
FSL addresses the problem of recognizing new categories from very few examples. This is directly applicable to detecting rare crash types (e.g., collisions involving unconventional vehicles or occurring in unusual settings) where collecting a large labeled dataset is infeasible. Prototypical networks [48], for example, learn a metric space where a new crash type can be classified based on its distance to a ”prototype” representation learned from just a handful of examples. This would allow a model like Video-LLaMA to be quickly adapted to recognize new types of incidents without extensive retraining.
VI-B3 A Proposed Hybrid Framework
A powerful future direction is to combine these approaches. A foundation model, such as a LLaMA-2 video model [31], could first be pre-trained via SSL on millions of hours of unlabeled public dashcam footage. This would create a robust general model of driving scenes. Then, this model could be specialized for crash detection via FSL, where it is fine-tuned on a very small number of labeled crash examples from datasets like DAD [38] or CADP [39]. This SSL-VAD approach, included conceptually in Table II, would be highly data-efficient and could achieve performance (e.g., >85absent85>85> 85% AUC on UCF-Crime) competitive with fully supervised methods while being far more scalable.
VI-C Discussion
The methods in Table II illustrate the shift from classical crash detection techniques [3, 6] to LLM-based systems, leveraging multimodal fusion and advanced reasoning as described in Section II-B and Section II-C. Early fusion (e.g., Video-LLaMA) and multimodal approaches (e.g., CrashLLM, VERA) excel in integrating visual and textual data, addressing occlusion and variability challenges. Training-free methods like LAVAD offer flexibility for real-time applications [37]. Performance metrics, such as CrashLLM’s 92% accuracy and VERA’s 86.55% AUC, highlight the superiority of LLM-based systems over traditional CNN-based methods, which struggle with complex scenarios [10].
This comparison reveals a clear trade-off in the current landscape. On one hand, methods like CrashLLM demonstrate state-of-the-art performance (¿92% accuracy) by leveraging custom-built multimodal architectures and large, domain-specific datasets like CrashEvent. However, this performance comes at the cost of significant computational resources and potential challenges in generalization. On the other hand, training-free approaches like LAVAD offer remarkable flexibility and efficiency, making them suitable for real-time deployment, but their performance (85.0% AUC) may not match specialized models in complex scenarios. Methods like VERA and Holmes-VAD strike a balance, focusing on explainability (a key challenge identified in Section VII) while achieving strong performance (86.55% AUC). This highlights a key decision point for ITS applications: is the primary goal maximum accuracy, real-time deployment, or human-interpretable output?
The inclusion of recent works like ScVLM [42], TrafficVLM [43], and CRASH [44] demonstrates the field’s rapid maturation. The focus is expanding from binary detection to a more nuanced understanding, including fine-grained conflict classification (ScVLM), dense narrative generation (TrafficVLM), and proactive anticipation (CRASH). This evolution is enabled by the availability of larger and more detailed datasets like SHRP 2 NDS [42] and the competitive environment of challenges like the AI City Challenge.
Datasets play a critical role, with CrashEvent, DAD, and UCF-Crime enabling robust training and evaluation (Section III). Evaluation metrics (Section V) like AUC, IoU, and BLEU provide a comprehensive assessment, but challenges remain in handling dataset imbalances and ensuring generalizability across diverse traffic conditions [11]. Future research should focus on developing hybrid fusion strategies, optimizing computational efficiency for real-time systems, and creating comprehensive, LLM-augmented datasets to further advance crash detection.
VII Challenges
The adoption of Large Language Models (LLMs) and Vision-Language Models (VLMs) for crash detection from video data has shown significant promise, as discussed in Sections VI and II-D. However, several challenges hinder their widespread deployment in intelligent transportation systems. This section outlines four key obstacles: data scarcity, multimodal alignment, reasoning and explainability, and real-time constraints. These challenges impact the robustness, interpretability, and practicality of LLM-based crash detection systems, necessitating further research to address them [11, 33].
VII-A Data Scarcity
LLM and VLM models require large, diverse datasets of labeled video-caption pairs to achieve robust performance in crash detection. However, such datasets are scarce, particularly for crash-specific scenarios, due to the rarity of crash events and the complexity of annotating video data [38]. While recent large-scale datasets like SHRP 2 NDS [42] have significantly improved data availability, they still represent a fraction of the data needed to cover all possible real-world scenarios. Existing datasets, such as the Dashcam Accident Dataset (DAD) and UCF-Crime, contain limited crash-related videos, often with coarse or incomplete annotations [38, 41]. For instance, DAD includes only 1,500 annotated videos, insufficient for training large-scale multimodal models like Video-LLaMA [27]. Synthetic datasets, such as those used in CrashLLM, attempt to address this by generating video-caption pairs, but they may lack the variability of real-world traffic conditions [33]. Data scarcity also exacerbates dataset imbalance, where non-crash videos dominate, leading to biased model predictions [11]. Efforts to create comprehensive, crash-specific datasets are critical to improving model generalization and performance [39].
VII-B Multimodal Alignment
Achieving temporal synchronization between video frames and language descriptions is a significant challenge, particularly in crash detection scenarios involving occlusions, low framerates, or complex traffic dynamics [29]. Multimodal alignment requires models to correlate visual events (e.g., a collision) with corresponding textual descriptions (e.g., “A car collided with a truck”), ensuring temporal and semantic consistency [27]. For example, frameworks like VERA struggle to align frame-level visual features with textual captions under occlusions, where key crash indicators may be obscured [29]. Low framerates, common in traffic surveillance videos, further complicate alignment by reducing temporal resolution, as noted in studies of classical optical flow methods [4]. Cross-attention mechanisms, as used in VAD-LLaMA, mitigate this by modeling long-range dependencies, but they increase computational complexity [34]. Developing efficient alignment techniques that handle occlusions and variable framerates remains an open research problem [11].
VII-C Reasoning and Explainability
While LLMs generate fluent crash reports, their outputs may be factually incorrect or inconsistent with visual inputs, a phenomenon known as hallucination [30]. This poses a critical challenge for crash detection, where accurate and interpretable outputs are essential for applications like autonomous driving and post-event analysis [33]. For instance, CrashLLM’s generative descriptions may inaccurately attribute crash causes (e.g., “sudden braking” instead of “lane change”) if visual cues are ambiguous [33]. Ensuring consistency between LLM-generated reports and visual evidence requires robust grounding techniques, such as those explored in VERA, which uses LLaVA-1.5 to verbalize crash rationales [29]. However, even advanced models struggle with causal reasoning under complex traffic scenarios, as noted in studies of multimodal forecasting [35]. Improving explainability through techniques like scene-graph grounding or causal intervention modules is a key open problem to enhance trust in LLM-based systems [26, 11].
VII-D Real-time Constraints
Real-time crash detection is critical for applications like accident response and autonomous vehicle decision-making, but LLM-based pipelines often exceed the computational limits required for real-time inference [37]. Large models, such as GPT-4 or Video-LLaMA, demand significant resources, with inference times ranging from hundreds of milliseconds to seconds per video clip, far exceeding the 30–100 ms latency required for real-time systems [30, 27]. For example, while the specific hardware is not detailed, the VADSK paper notes that the foundational models used within Holmes-VAD have inference speeds ranging from approximately 2 to 6 seconds per frame, clearly outside the real-time window for immediate response [58]. This highlights a significant performance gap that needs to be addressed for practical deployment.
Optimizing model architectures (e.g., using lightweight adapters like LoRA [32]), developing efficient temporal modules, and leveraging quantization techniques are essential to meet real-time constraints while maintaining performance [11]. For example, Holmes-VAD’s multi-modal LLM achieves high accuracy but incurs high computational costs, limiting its use in resource-constrained environments [56]. Training-free approaches like LAVAD reduce latency by eliminating fine-tuning, achieving 85.0% AUC with lower overhead, but they may sacrifice precision in complex scenarios [37]. Classical methods, such as frame differencing, offer lower latency but lack the robustness of LLMs [3]. A quantitative comparison of inference times and hardware requirements across models remains a challenge due to a lack of standardized reporting in the literature, but the available data suggests a clear trade-off between model complexity, accuracy, and processing speed.
VII-E Ethical and Fairness Considerations
Beyond technical hurdles, the deployment of LLM-based crash detection systems raises significant ethical and fairness concerns. Datasets such as DAD and UCF-Crime, while foundational, may contain inherent biases by over-representing specific geographic regions, weather conditions, or vehicle types. Models trained on such data may perform poorly in underrepresented scenarios, leading to inequitable safety outcomes. Furthermore, the use of dashcam and surveillance footage introduces profound privacy risks, as this data can capture identifiable information like faces and license plates. The societal impact of large-scale video data collection and analysis requires careful consideration of data anonymization techniques and transparent policies regarding data usage and storage.
Algorithmic fairness is another critical issue. Biased data can lead to biased models, where false positive or negative rates are disproportionately higher for certain demographic groups or environments. While methods like Holmes-VAD aim to address unbiased detection, the risk remains that automated systems could perpetuate or even amplify existing societal biases, with serious consequences for insurance liability and legal accountability. Therefore, developing fair, transparent, and privacy-preserving systems is not just a technical challenge but an ethical imperative for the responsible advancement of this technology.
VII-F Robustness and Generalization Analysis
A critical challenge for LLM-based crash detection models is their ability to generalize to scenarios not well represented in the training data. While models may perform well on benchmarks, their robustness in unseen or out-of-distribution (OOD) scenarios—such as rural roads, adverse weather (e.g., snow, heavy rain), or different geographic regions—remains a significant concern. This limitation poses a risk of overfitting to the visual and contextual cues prevalent in the training data, leading to a sharp degradation in performance when deployed in the real world [59].
VII-F1 Domain Generalization
To address this, domain generalization techniques are essential. These methods aim to learn domain-invariant features that are robust to shifts in data distribution. Techniques such as domain-invariant feature learning, meta-learning, and adversarial domain adaptation can help models generalize from multiple source domains to unseen target domains [60]. Data augmentation, particularly using high-fidelity simulators like CARLA, allows for the creation of diverse training data by simulating varied weather and lighting conditions, which can significantly improve model robustness [61].
VII-F2 Out-of-Distribution (OOD) Detection
Complementary to domain generalization is Out-of-Distribution (OOD) detection, which focuses on identifying whether a given input is from a novel, unseen distribution. In crash detection, this is crucial for flagging rare events (e.g., unusual multi-vehicle pile-ups, crashes involving non-standard vehicles) that the model was not trained to recognize. By equipping models like LAVAD or VERA with an OOD detection mechanism, the system could defer to a human operator or a fail-safe mode when it encounters a scenario it cannot confidently analyze, thereby increasing safety. Methods based on uncertainty estimation are particularly promising for this task, as they can quantify the model’s confidence in its predictions [62].
VII-F3 Cross-Dataset Evaluation and Quantitative Analysis
A rigorous evaluation methodology is necessary to properly assess generalization. This should involve comprehensive cross-dataset evaluation, where models trained on one dataset (e.g., DAD) are tested on another with different characteristics (e.g., UCF-Crime or the more diverse SHRP 2 NDS) [40, 42]. A quantitative analysis of this performance drop, for example, by reporting the decrease in AUC or F1-score, provides a clear measure of a model’s generalization capability. Future work should standardize this practice and ideally present results in a comparative table, analyzing performance drops with statistical significance tests to build a clearer picture of which models are most robust.
VII-F4 Augmentation with Synthetic Data
The scarcity of diverse, real-world crash data is a major bottleneck. Synthetic data generation offers a scalable solution. Simulators like CARLA and SUMO can be used to generate vast amounts of labeled data covering a wide range of scenarios, including edge cases that are rare in the real world [61, 59]. By training or fine-tuning models on a combination of real and synthetic data, it is possible to significantly improve their ability to generalize to new environments and situations, making them more reliable for real-world deployment.
The challenges of data scarcity, multimodal alignment, reasoning and explainability, real-time constraints, and generalization underscore the complexities of deploying LLM and VLM-based crash detection systems. Data scarcity limits model generalization, necessitating larger, more diverse datasets [39]. Multimodal alignment issues highlight the need for advanced fusion strategies, such as cross-attention, to handle occlusions and low framerates [34]. Reasoning and explainability challenges demand robust grounding techniques to ensure factual accuracy [29]. Real-time constraints require optimized architectures to balance latency and performance [37]. Finally, ensuring robust generalization requires a combination of advanced techniques like domain generalization, OOD detection, and the use of synthetic data, all validated through rigorous cross-dataset evaluation. Addressing these challenges will enhance the reliability and applicability of LLM-based crash detection, building on the advancements discussed in Section VI and paving the way for future research in intelligent transportation systems.
VIII Ethical, Legal, and Societal Implications
Beyond the technical challenges, the deployment of LLM-based crash detection systems carries profound ethical, legal, and societal implications that demand careful consideration. As these technologies become more integrated into safety-critical domains, a proactive approach to addressing their broader impact is essential for responsible innovation and public trust.
VIII-A Bias and Fairness
A primary concern is the risk of algorithmic bias. Datasets used to train crash detection models, such as DAD and UCF-Crime, often over-represent urban environments and specific weather conditions, potentially leading to biased models that underperform in rural or atypical scenarios [63]. This can result in inequitable safety outcomes, where the system is less reliable for certain communities. To mitigate this, researchers propose debiasing techniques like reweighting underrepresented data samples and adversarial training to create more robust and equitable models. Ensuring fairness requires a concerted effort to build diverse and representative datasets that reflect the full spectrum of real-world driving conditions.
VIII-B Privacy Risks
The use of dashcam and surveillance footage for crash detection raises significant privacy concerns, as this data can capture personally identifiable information such as faces and license plates. The large-scale collection and analysis of such video data require robust privacy-preserving measures. Techniques like homomorphic encryption, which allows computation on encrypted data, and differential privacy, which adds statistical noise to data to protect individual identities, are promising solutions [64]. Furthermore, strict adherence to data protection regulations like GDPR and CCPA is necessary to ensure that data is handled responsibly and transparently.
VIII-C Legal Accountability
The increasing autonomy of crash detection systems introduces complex questions of legal accountability. When an LLM-based system like CrashLLM or VERA generates a crash report that is used in insurance claims or legal proceedings, determining liability in the event of an error becomes challenging. Is the responsibility with the vehicle owner, the software developer, or the vehicle manufacturer? A clear legal framework is needed to address these new liability scenarios. Recent surveys on AI safety assurance emphasize the need for new standards and regulations that cover the entire lifecycle of AI systems, from design and testing to deployment and maintenance, to establish clear lines of accountability [65].
VIII-D Societal Impact and Public Trust
The societal impact of LLM-based crash detection includes significant benefits, such as the potential for faster emergency response and a reduction in road fatalities. However, there are also risks, including over-reliance on automated systems and public apprehension. Building public trust is paramount for the successful adoption of these technologies. Surveys on public trust in AI reveal that while citizens see the potential benefits, they also harbor significant concerns about job loss, privacy, and misuse of the technology by authorities [66]. Fostering trust requires transparent communication about how the technology works, its limitations, and the safeguards in place. Engaging the public in the development and oversight process and ensuring that the systems are explainable and accountable will be key to building public confidence [63].
VIII-E Case Study: A Hypothetical Scenario
Consider a hypothetical scenario where an LLM-based crash detection system is widely deployed in a country. The system works exceptionally well in urban areas but has not been sufficiently trained on data from rural regions with unpaved roads and different types of agricultural vehicles. An accident occurs on a rural road involving a tractor, but the system fails to detect it in a timely manner, delaying emergency services. The legal investigation reveals the system’s biased performance. This scenario highlights the severe consequences of dataset bias, leading to legal challenges against the manufacturer for negligence, regulatory fines for deploying an unsafe product, and a significant erosion of public trust in AI-powered safety systems. It underscores the critical need for comprehensive testing, fair representation in training data, and robust regulatory oversight before widespread deployment.
IX Integration with Autonomous Driving Ecosystems
The practical value of LLM-based crash detection is maximized when it is seamlessly integrated into the broader autonomous vehicle (AV) ecosystem. This integration enhances the vehicle’s perception-planning-action loop, providing critical, real-time insights that can prevent accidents or mitigate their severity. This section explores how these models interface with core AV modules.
IX-A AV Pipeline Integration
A standard AV software architecture follows a perception-planning-action pipeline [67]. LLM-based crash detection serves as an advanced component of the perception stack. The outputs from a model like VERA or CrashLLM—such as a textual description (”imminent rear-end collision due to sudden braking”) or a classification of risk—can be fed directly into the AV’s decision-making or planning module. This allows the AV to move beyond simple object detection and react to a nuanced understanding of the scene’s dynamics. For example, a high-risk prediction could trigger an emergency braking maneuver or a swerving action, demonstrating a direct link between semantic crash understanding and vehicle control.
IX-B Multi-Sensor Fusion
While cameras are the primary input for the VLM-based methods discussed, their fusion with other sensor modalities like LiDAR and radar is crucial for building a robust perception system for AVs. LiDAR provides precise 3D geometric data, which is less affected by adverse weather, while radar offers reliable velocity information. A multi-sensor fusion approach would combine the semantic, contextual understanding from an LLM/VLM with the geometric and velocity data from LiDAR/radar [68]. For example, a VLM might identify a ”pedestrian looking at their phone,” and this semantic flag could be fused with LiDAR data confirming their trajectory towards the road, leading to a more reliable and timely pre-emptive braking action than either sensor could achieve alone. Extending frameworks like Video-LLaMA to incorporate these additional data streams is a promising research direction.
IX-C Real-time Performance and Latency
A critical requirement for any system integrated into an AV pipeline is real-time performance, typically with a latency of under 100 ms, and often as low as 30 ms for critical safety functions. As discussed in Section VII, large models present a significant challenge. To address this, techniques such as model quantization, which reduces the precision of the model’s weights, can significantly decrease model size and speed up inference with minimal loss in accuracy [69]. Lightweight models, such as MiniGPT-4 [45], are specifically designed for resource-constrained environments and offer a viable path for deploying advanced reasoning capabilities on embedded AV hardware.
IX-D Crash Anticipation as a Proactive Safety Measure
Modern AV safety is shifting from reactive collision avoidance to proactive crash anticipation. Models like CRASH [44], which are trained to predict the likelihood of a crash several seconds before it occurs, are perfectly aligned with this paradigm. The output of such an anticipation model can serve as a critical input to the AV’s planning module, allowing it to take evasive action—such as reducing speed or changing lanes—well before a dangerous situation becomes unavoidable. This proactive stance, informed by the rich contextual understanding of video data, represents a significant step towards achieving Level 4 and Level 5 autonomy.
X Quantitative Benchmarking and Ablation Studies
To provide a clearer, more rigorous comparison of the methods discussed, this section proposes a framework for unified quantitative benchmarking and outlines the need for systematic ablation studies.
X-A Cross-Dataset Generalization
A crucial test of robustness is cross-dataset generalization. Models should be trained on one dataset (e.g., CADP) and evaluated on another (e.g., BDD100K) without fine-tuning. This tests the model’s ability to generalize to new environments, camera types, and traffic patterns. Reporting the performance drop in such scenarios would provide a quantitative measure of a model’s generalization capabilities, which is currently a major gap in the literature.
X-B Ablation Studies
To understand the contribution of individual architectural components, systematic ablation studies are needed. For instance, in a model like VERA, one could ablate the language-based reasoning module to quantify its impact on performance compared to using only the visual features. Similarly, for Video-LLaMA, ablating the temporal module would reveal how much the model relies on temporal cues versus single-frame analysis. Such studies are essential for understanding what makes these models effective and for guiding future research [70].
X-C Computational Metrics
For practical deployment, especially in AVs, computational performance is as important as accuracy. Future benchmarking should include a systematic comparison of inference latency (in milliseconds) and computational requirements (e.g., GPU memory, TFLOPs) on standardized hardware. For example, comparing the performance of a heavyweight model like Holmes-VAD against a lightweight, training-free model like LAVAD would highlight the trade-offs between accuracy and efficiency, providing critical insights for real-world applications where resources are constrained [58].
XI Future Directions
The challenges outlined in Section VII and the advancements in LLM-based crash detection methods discussed in Section VI highlight the need for innovative approaches to enhance the robustness, scalability, and applicability of these systems in intelligent transportation systems. This section proposes key future directions to address current limitations and advance the field.
XI-A Synthetic Training Data
The scarcity of labeled video-caption pairs for crash detection, as noted in Section VII, limits the generalization of LLMs and VLMs [38, 39]. To address this, synthetic training data generated through simulation platforms like CARLA or SUMO offers a promising solution [40]. These platforms can create diverse crash scenarios, including rare events (e.g., multi-vehicle collisions under adverse weather), paired with detailed textual narratives (e.g., “A truck skidded on a wet road, colliding with a sedan”). For instance, CARLA’s high-fidelity simulations have been used to generate video data for autonomous driving tasks, which can be extended to crash-specific datasets with annotations for temporal events and causal factors [40]. Synthetic data can augment existing datasets like DAD and UCF-Crime, reducing dataset imbalance and improving model robustness [41]. Recent work on CrashLLM demonstrates the potential of synthetic video-caption pairs, though real-world variability remains a challenge [33]. Future efforts should focus on hybrid datasets combining synthetic and real-world data, leveraging techniques like domain adaptation to ensure transferability [23]. Such approaches can mitigate data scarcity, enabling models to handle diverse traffic conditions and rare crash scenarios [11].
XI-B Video-grounded QA Benchmarks
Current crash detection systems, such as VERA and Holmes-VAD, excel in generating descriptive crash reports but lack robust question-answering (QA) capabilities for interactive analysis [29, 56]. Developing video-grounded QA benchmarks specific to crash scenarios can address this gap, enhancing explainability and supporting applications like post-event forensics and autonomous vehicle decision-making [35]. These benchmarks would include datasets with video clips paired with questions and answers about crash details (e.g., “What caused the collision?” or “Which vehicle was at fault?”), building on frameworks from recent literature. For example, a QA dataset could extend the CADP dataset with annotated questions derived from crash reports, enabling models to reason about temporal and causal relationships [39]. Recent advances in video-language models, such as Video-LLaMA, provide a foundation for processing multimodal inputs, but crash-specific QA datasets are scarce [27]. Evaluation metrics like BLEU, METEOR, and CIDEr, used in captioning tasks, can be adapted to assess QA performance, ensuring factual accuracy [53, 54, 55]. Creating such benchmarks will improve model interpretability and enable interactive crash analysis, addressing the reasoning challenges noted in Section VII [30].
XI-C Fine-tuned VLMs
While pre-trained VLMs like BLIP-2 and Flamingo offer strong generalization, their performance in crash-specific scenarios is limited by the lack of domain-specific fine-tuning [25, 26]. Fine-tuning VLMs on crash-specific scenes and annotations can enhance their ability to detect and describe complex crash events, addressing multimodal alignment issues [29]. For instance, fine-tuning Video-LLaMA on datasets like DAD or BDD100K, with annotations for crash types, severity, and temporal dynamics, can improve temporal synchronization and feature extraction [27, 40]. Techniques like LoRA (Low-Rank Adaptation) enable efficient fine-tuning of large models, reducing computational overhead while adapting to crash-specific contexts [32]. Recent work on VERA demonstrates the benefits of fine-tuning LLaVA-1.5 for explainable crash detection, achieving high AUC and BLEU scores [29]. Future efforts should focus on curating crash-specific annotations, incorporating temporal models like TimeSformer or SlowFast networks to capture dynamic events [52, 46]. Such fine-tuned VLMs will enhance accuracy and robustness, particularly in challenging scenarios involving occlusions or low framerates, as discussed in Section VII [11].
XI-D Multilingual and Low-resource Models
Most LLM and VLM-based crash detection systems rely on English-language datasets and descriptions, limiting their applicability in non-English-speaking regions [33]. Expanding to multilingual and low-resource models can improve accessibility, enabling crash detection systems to process descriptions in languages like Spanish, Mandarin, or Hindi. Multilingual pre-training, as explored in models like MiniGPT-4, provides a foundation for handling diverse linguistic inputs [45]. For example, fine-tuning Video-LLaMA on multilingual crash datasets, such as those derived from global traffic surveillance, can enable cross-linguistic crash analysis [27]. Low-resource models, optimized for environments with limited computational resources, are equally critical for deployment in developing regions. Techniques like knowledge distillation or training-free approaches, as used in LAVAD, can reduce model complexity while maintaining performance [37]. Initiatives like TrafficLens demonstrate the potential of lightweight models for video-to-text analysis in resource-constrained settings [36]. Future work should prioritize multilingual dataset creation and efficient model architectures, ensuring equitable access to crash detection technologies across diverse linguistic and economic contexts [11, 10].
XI-E Edge Computing and Distributed Learning
The real-time and privacy-sensitive nature of crash detection necessitates a shift towards edge computing and distributed learning architectures. These approaches address the latency and data privacy challenges inherent in cloud-centric models.
XI-E1 Edge-Optimized Models
Deploying complex VLMs on resource-constrained edge devices (e.g., in-vehicle computers) requires significant model optimization. Techniques like pruning (removing redundant model weights), quantization (reducing the numerical precision of weights), and knowledge distillation (training a smaller model to mimic a larger one) are crucial for reducing model size and latency [69]. Research in edge-based video analytics for ITS has shown that such optimizations can enable real-time performance on embedded hardware [71]. For example, a quantized and pruned version of a model like MiniGPT-4 could be deployed on edge cameras for real-time crash detection, as its smaller footprint would meet the strict computational budget of in-vehicle systems.
XI-E2 Federated Learning for Privacy-Preserving Training
Federated learning offers a powerful paradigm for training models on decentralized data sources, such as data from a fleet of vehicles or a network of traffic cameras, without compromising privacy. In this approach, the model is trained locally on each device, and only the model updates (gradients), not the raw data, are sent to a central server for aggregation [72]. This is particularly valuable for crash detection, as it allows a global model to learn from a diverse range of driving scenarios and environments without requiring sensitive video data to be uploaded to the cloud, thus preserving the privacy of drivers.
XI-E3 Distributed Inference for Low Latency
For very large models, even after optimization, inference may be too slow for a single edge device. Distributed inference, where the computational load is split between the edge and the cloud, offers a solution. For example, an initial, low-latency analysis could be performed on the edge device to detect potential anomalies, and if a high-risk event is flagged, more detailed video data could be sent to a powerful cloud-based model for a more thorough analysis and descriptive summary [73]. This hybrid approach balances the need for low latency with the desire for deep, semantic understanding of crash events. A case study could involve deploying a lightweight variant of LAVAD on edge cameras to continuously monitor for anomalies, with its outputs triggering a more powerful VLM in the cloud for detailed causal analysis, with latency being a key evaluation metric on a dataset like CCD [37, 44].
XI-F Active Learning and Human-in-the-Loop (HITL)
To further address data scarcity and improve model reliability, active learning and Human-in-the-Loop (HITL) systems present a promising frontier. These approaches can create a virtuous cycle of data collection, annotation, and model refinement, making the entire system more efficient and trustworthy.
XI-F1 Active Learning for Efficient Annotation
Active learning can significantly reduce the manual annotation burden by intelligently selecting the most informative data for labeling [74]. For crash detection, an active learning system could prioritize video clips where the model is most uncertain about its predictions [75]. For example, after an initial model is trained on a small subset of DAD or CADP, it could be used to score a large, unlabeled pool of traffic videos. The videos with the highest uncertainty would then be sent to human annotators. This targeted annotation strategy ensures that labeling efforts are focused on the most challenging and valuable examples, accelerating model improvement and reducing the cost of creating large-scale datasets like DAD or CADP [38, 39].
XI-F2 Human-in-the-Loop for Validation and Refinement
HITL systems are essential for safety-critical applications, where the cost of an error is high. In the context of crash detection, a HITL system would involve human experts validating the outputs of models like VERA or CrashLLM [76]. This is particularly important for correcting ”hallucinations” or factual inaccuracies in LLM-generated explanations of crash events. By providing corrective feedback, human experts can iteratively refine the model’s performance and ensure its outputs are reliable and trustworthy. This process also generates a valuable dataset of human-verified crash descriptions, which can be used for further fine-tuning. Evaluation metrics for such a system could go beyond standard captioning scores like BLEU to include human agreement scores and measures of annotation efficiency, providing a more holistic assessment of the system’s performance.
The proposed future directions—synthetic training data, video-grounded QA benchmarks, fine-tuned VLMs, and multilingual and low-resource models—address the core challenges of LLM-based crash detection systems. Synthetic data generation can mitigate data scarcity, building on datasets like BDD100K and CADP [40, 39]. QA benchmarks will enhance explainability, complementing methods like VERA [29]. Fine-tuned VLMs will improve multimodal alignment, leveraging architectures like TimeSformer [52]. Multilingual and low-resource models will ensure global applicability, aligning with real-time needs [37]. These directions, supported by foundational work in vision-language pre-training [23, 24, 28] and temporal modeling [51, 49], pave the way for robust, interpretable, and accessible crash detection systems, advancing the field beyond the methods discussed in Section VI [11].
XII Conclusion
The advent of Large Language Models (LLMs) and Vision-Language Models (VLMs) has transformed crash detection from video data, enabling high-level reasoning that bridges visual observations with linguistic descriptions. This survey comprehensively reviewed advancements in LLM-based crash detection from 2023 to 2024, covering models, datasets, taxonomies, challenges, and future directions, with the goal of guiding researchers toward building reliable and explainable multimodal systems for safety-critical applications in intelligent transportation systems [11, 33]. By synthesizing recent developments, this work underscores the potential of LLMs to revolutionize traffic safety while identifying critical areas for further innovation.
Section VI analyzed state-of-the-art methods, such as CrashLLM, VERA, Holmes-VAD, and Video-LLaMA, which leverage multimodal fusion strategies (e.g., early, late, cross-attention) to achieve high accuracy and explainability [33, 29, 56, 27]. These methods outperformed classical approaches, such as CNN-based systems, by integrating visual and textual data to address complex scenarios involving occlusions and variable lighting [3, 6]. For instance, CrashLLM’s generative capabilities produce detailed crash descriptions, achieving over 92% accuracy, while VERA’s verbalized learning enhances post-event analysis with 86.55% AUC [33, 29]. The survey also introduced a taxonomy of fusion strategies and LLM roles (e.g., passive, generative), providing a structured framework for understanding these advancements [11, 27].
Key datasets, including BDD100K, CADP, and UCF-Crime, have been instrumental in training these models, though their limitations—such as limited crash-specific videos and coarse annotations—pose challenges [40, 39, 41]. The recent emergence of large-scale, finely annotated datasets like SHRP 2 NDS [42] is a critical step forward. Section VII detailed these obstacles, highlighting data scarcity, multimodal alignment difficulties, reasoning errors (e.g., hallucination), and real-time computational constraints [38, 30, 37]. For example, the Dashcam Accident Dataset (DAD) contains only 1,500 videos, insufficient for large-scale VLM training, while real-time systems like Video-LLaMA struggle with latency requirements of 30–100 ms [38, 27]. These challenges underscore the need for innovative solutions to enhance robustness and practicality.
Section XI proposed four directions to address these issues: generating synthetic training data using simulators like CARLA, developing video-grounded QA benchmarks, fine-tuning VLMs for crash-specific scenarios, and creating multilingual and low-resource models [40, 32, 45]. These strategies aim to mitigate data scarcity, improve temporal alignment, enhance explainability, and ensure global accessibility. For instance, synthetic data can augment datasets like CADP, while QA benchmarks can enable interactive crash analysis [39]. Fine-tuning with techniques like LoRA and multilingual pre-training with models like MiniGPT-4 can further advance performance and equity [32, 45].
The implications of these advancements are profound for safety-critical applications, such as autonomous driving, traffic surveillance, and accident response. Methods like Holmes-VAD and LAVAD demonstrate the potential for unbiased and training-free crash detection, while TrafficVLM highlights the power of dense captioning for post-event analysis [56, 37, 43]. Foundational models, including CLIP, Flamingo, and GPT-4, have laid the groundwork for robust feature extraction and reasoning, supported by evaluation metrics like BLEU, METEOR, and CIDEr [23, 26, 30, 53, 54, 55]. Temporal modeling architectures, such as TimeSformer, SlowFast, and ConvLSTM, further enhance video understanding, addressing dynamic crash scenarios [52, 46, 51].
This survey encourages researchers to build on these foundations, leveraging insights from recent methods and proposed directions to develop multimodal systems that are accurate, interpretable, and deployable in real-world settings. By addressing challenges through innovations like synthetic data and fine-tuned VLMs, the field can move toward standardized benchmarks and efficient architectures, as seen in works like CLIP4Clip and scaling approaches [28, 24]. Ultimately, we hope this survey serves as a catalyst for advancing LLM-based crash detection, fostering safer and more intelligent transportation systems [11, 10].

### v2

@@
 Figure 1: Taxonomy of LLM-based crash detection in videos categorized by fusion strategy, LLM role, and input granularity.
+LatencyAccuracy100500100015002000+80859095Deployment ReadyConditionalResearch OnlyCrashLLMVERAVideo-LLaMALAVADHolmes-VADCRASHIDEAL ZONE
+Figure 2: Accuracy vs.
+Latency Trade-off Analysis for LLM-based Crash Detection Methods.
+The green zone represents deployment-ready systems (<500<500 ms), yellow shows conditional deployment scenarios, and red indicates research-only systems.
 I Introduction
-Crash detection is vital for autonomous vehicles, traffic monitoring, and emergency response.
-Traditionally, it has relied on heuristics or vision-based CNNs using motion cues.
-However, these methods struggle with generalization, occlusion, and complex temporal dependencies.
-Large Language Models (LLMs), such as GPT-4 and PaLM, have demonstrated impressive reasoning across modalities, especially when integrated with visual encoders like CLIP, Flamingo, or BLIP.
-This integration introduces a new paradigm: interpreting crash-related events in video via language-guided reasoning, captioning, and event prediction.
-This survey aims to provide a structured overview of recent approaches leveraging LLMs/VLMs for crash understanding in video, summarizing models, datasets, and evaluation strategies.
+Crash detection from video data has been a critical challenge in intelligent transportation systems for over two decades.
+Early approaches, dating back to the early 2000s, relied on classical computer vision techniques such as optical flow analysis, background subtraction, and trajectory-based anomaly detection [1, 2, 3].
+These methods, while computationally efficient and interpretable, achieved limited accuracy (70-85%) and struggled with fundamental challenges including environmental sensitivity, occlusion handling, and poor generalization across different traffic scenarios [4, 5].
+The deep learning revolution (2015-2020) brought significant improvements, with CNN-based approaches and specialized architectures like Two-Stream networks, 3D CNNs, and SlowFast networks achieving 85-95% accuracy on video analysis benchmarks [6, 7, 8, 9].
+Video anomaly detection methods, particularly those developed for surveillance applications, provided direct foundations for crash detection through weakly-supervised learning and future frame prediction approaches [10, 11, 12].
+However, these methods remained limited by their black-box nature, substantial data requirements, and inability to provide interpretable explanations for safety-critical decisions.
+The emergence of multimodal learning (2020-2022) began to address some of these limitations through vision-language models that could incorporate semantic understanding alongside visual processing [13, 14, 15, 16].
+These approaches demonstrated the potential for cross-modal reasoning and zero-shot generalization, though they remained constrained by limited temporal modeling capabilities and basic language understanding.
+Large Language Models (LLMs), such as GPT-4 and PaLM, have recently demonstrated impressive reasoning capabilities across modalities, especially when integrated with visual encoders like CLIP, Flamingo, or BLIP [17, 18, 19].
+This integration introduces a transformative paradigm: interpreting crash-related events in video through language-guided reasoning, detailed captioning, causal analysis, and predictive event modeling.
+Unlike previous approaches that focused primarily on detection accuracy, LLM-based systems can provide rich explanations (e.g., “A vehicle swerved due to sudden braking, causing a rear-end collision”), perform causal reasoning about crash sequences, and adapt to new incident types through few-shot learning.
+However, as Table I illustrates, this evolution has come with new challenges.
+While LLM-based approaches offer unprecedented capabilities in contextual understanding and explanation generation, they face significant deployment barriers including high computational costs (200ms-2s latency), hallucination risks, and substantial memory requirements that limit real-world applicability.
+This survey provides the first comprehensive analysis of this fundamental paradigm shift, examining how the field has evolved from reactive, pixel-based anomaly detection to proactive, context-aware event interpretation powered by language-guided reasoning.
+We present a unified taxonomy of LLM-based approaches, analyze their performance against classical and deep learning baselines, and critically assess their readiness for real-world deployment.
+Our analysis reveals both the transformative potential and the significant engineering challenges that must be addressed to realize practical LLM-based crash detection systems.
+TABLE I: Evolution of Crash Detection Methods: Performance and Limitations Across Different Eras
+Era Period Typical Accuracy Latency Key Methods Main Advantages Critical Limitations Classical 2000-2015 70-85% <<100ms Optical flow, background subtraction, trajectory analysis, SVM Fast processing, interpretable rules, low computational cost Manual feature engineering, environmental sensitivity, poor generalization Deep Learning 2015-2020 85-95% 100-500ms CNN, 3D CNN, Two-Stream, I3D, SlowFast, LSTM Automatic feature learning, better generalization, robust to variations Black-box nature, data hungry, limited context understanding Early Multimodal 2020-2022 88-92% 300-800ms CLIP, UNITER, VideoCLIP, BLIP Cross-modal understanding, zero-shot capabilities, semantic reasoning Limited temporal modeling, basic language understanding LLM-based 2023-2025 85-92% 200ms-2s Video-LLaMA, VERA, CrashLLM, Holmes-VAD, HybridMamba Rich explanations, causal reasoning, contextual understanding, few-shot learning High computational cost, hallucination risk, deployment challenges
 II Background and Related Work
-The domain of video-based crash detection has evolved significantly, moving from classical computer vision techniques to modern deep learning and language-guided approaches.
-This section provides background on the foundational methods that preceded LLM integration and reviews the key developments in Vision-Language Models (VLMs) and their application to video understanding, which together set the stage for the methods surveyed in this paper.
-II-A Crash Detection in Video
-Classical approaches to crash detection in video include frame differencing, optical flow, background subtraction, and CNN-based classification.
-These methods rely heavily on low-level motion or scene cues, such as changes in pixel intensity, motion vectors, or spatial patterns, to identify anomalies indicative of crashes.
-Early research in this area often focused on trajectory-based analysis, where anomalies were detected if objects deviated from learned normal paths [1, 2].
-However, these methods were sensitive to occlusion and relied on continuous tracking.
-Subsequent approaches shifted to using low-level features that were more robust.
-Frame differencing detects sudden changes between consecutive frames, often used to identify abrupt vehicle movements [3].
-Optical flow analyzes motion patterns to track vehicle trajectories, enabling detection of collisions or erratic behavior [4, 5].
-Background subtraction isolates moving objects from static scenes, useful for detecting crash-related anomalies in traffic videos [6].
-Other traditional methods leveraged hand-crafted features like Histograms of Oriented Gradients (HOG) and Histograms of Optical Flow (HOF) to capture appearance and motion information [7].
-While effective in controlled settings, these approaches often struggle with the ambiguity and variability of real-world driving scenarios, such as varying lighting conditions or complex traffic interactions [8, 9].
-This led to the adoption of deep learning, with CNN-based classification, such as using ResNet or YOLO, leveraging deep learning to classify video frames or regions as crash or non-crash events, often combining spatial and temporal features through architectures like Convolutional Long Short-Term Memory (Conv-LSTM) [3, 10].
-However, even these methods can be limited, prompting the exploration of more robust methods like Large Language Models (LLMs) and Vision-Language Models (VLMs) for enhanced crash detection [11].
+Video-based crash detection has evolved through distinct phases: from classical computer vision techniques (2000-2015) to deep learning approaches (2015-2020), and finally to modern language-guided methods (2020-present).
+This section provides a comprehensive review of this evolution, establishing the foundational context necessary to understand current LLM-based approaches and their advantages over existing methods.
+II-A Classical Crash Detection Methods
+Early approaches to crash detection relied primarily on low-level image processing techniques and traditional machine learning methods.
+These foundational works established the core challenges and evaluation frameworks that continue to influence modern research.
+II-A1 Motion-Based Detection
+Classical motion-based approaches formed the backbone of early crash detection systems, leveraging fundamental computer vision techniques to identify anomalous vehicle behavior.
+Optical Flow and Frame Differencing: Early systems used optical flow analysis to track vehicle trajectories and detect sudden changes indicative of crashes [1].
+Kamijo et al.
+pioneered intersection monitoring using optical flow to identify collision events, achieving 85% accuracy on controlled datasets.
+Frame differencing techniques detected abrupt changes between consecutive frames, with Morris and Trivedi providing a comprehensive survey of trajectory-based analysis methods [2].
+However, these approaches struggled with occlusions, varying lighting conditions, and camera jitter, limiting their real-world applicability.
+Background Subtraction: Background subtraction methods isolated moving objects from static scenes to identify crash-related anomalies [3].
+Piciarelli et al.
+developed trajectory-based anomaly detection using statistical models of normal traffic flow, achieving detection rates of 78% on highway surveillance data.
+While effective in controlled environments, these methods suffered from sensitivity to environmental changes and required extensive parameter tuning for different scenarios.
+Trajectory Analysis: Trajectory-based methods analyzed vehicle paths to identify deviations from normal traffic patterns.
+Early work by Jiang et al.
+used spatiotemporal context to detect anomalous events, incorporating both spatial and temporal features for improved robustness [20].
+Calderara et al.
+employed spectral graph analysis to model normal trajectory patterns, detecting crashes as significant deviations from learned behaviors [21].
+These approaches achieved 70-80% accuracy on controlled datasets but struggled with complex multi-vehicle scenarios.
+II-A2 Traditional Machine Learning Approaches
+As computational power increased, researchers began applying traditional machine learning techniques to crash detection, moving beyond simple heuristics to data-driven approaches.
+Support Vector Machines (SVMs): Li et al.
+pioneered the use of SVMs for crash injury severity classification, demonstrating superior performance compared to traditional statistical methods [22].
+Their approach achieved 82% accuracy on real-world crash data, establishing SVMs as a viable tool for crash-related prediction tasks.
+However, SVMs required careful feature engineering and struggled with temporal dependencies in video data.
+Neural Networks: Early neural network approaches, such as those developed by Abdelwahab and Abdel-Aty, used artificial neural networks to predict driver injury severity at signalized intersections [4].
+These methods achieved 76% accuracy and demonstrated the potential of learning-based approaches, though they were limited by computational constraints and small datasets.
+Accident Detection Systems: Ki and Lee developed comprehensive accident detection systems using image processing combined with traditional pattern recognition techniques [23].
+Their system integrated multiple visual cues including motion patterns, shape analysis, and temporal consistency, achieving 83% detection accuracy on controlled test scenarios.
+However, the system required manual parameter tuning and failed to generalize across different traffic scenarios.
+II-A3 Performance Analysis and Limitations
+Classical approaches established important benchmarks and revealed fundamental challenges that continue to influence modern research.
+Performance Characteristics: Classical methods typically achieved 70-85% accuracy on controlled datasets, with significant performance degradation in real-world scenarios.
+Motion-based approaches excelled in detecting obvious crashes with clear trajectory changes but struggled with subtle incidents or partially occluded events.
+Traditional ML methods showed better generalization but required extensive feature engineering and domain expertise.
+Fundamental Limitations: Several critical limitations motivated the transition to deep learning approaches:
+  • Feature Engineering Bottleneck: Manual feature design required extensive domain knowledge and failed to capture complex spatiotemporal patterns.
+  • Environmental Sensitivity: Performance degraded significantly under varying lighting conditions, weather changes, or camera viewpoints.
+  • Occlusion Handling: Most methods failed when key visual information was occluded by other vehicles or objects.
+  • Temporal Modeling: Limited ability to model long-term temporal dependencies crucial for understanding crash sequences.
+  • Generalization: Poor transfer performance across different geographic regions, traffic patterns, or camera setups.
+These limitations directly motivated the development of deep learning approaches, which promised to learn features automatically and handle complex spatiotemporal patterns more effectively.
+The transition from classical to deep learning approaches marked a paradigm shift in video analysis, moving from hand-crafted features to learned representations.
+This evolution was particularly impactful for crash detection, where complex spatiotemporal patterns required more sophisticated modeling approaches.
+II-A4 CNN-Based Video Understanding (2014-2018)
+The introduction of convolutional neural networks to video analysis revolutionized the field, providing automatic feature learning capabilities that surpassed hand-crafted approaches.
+Two-Stream Networks: Simonyan and Zisserman introduced the two-stream architecture, processing RGB frames and optical flow separately before fusion [6].
+This approach achieved 88% accuracy on action recognition benchmarks, demonstrating the value of explicitly modeling motion alongside appearance.
+For crash detection, two-stream networks enabled better understanding of both vehicle appearance and motion dynamics, though they required pre-computed optical flow.
+3D Convolutional Networks: Tran et al.
+pioneered 3D CNNs for spatiotemporal feature learning, extending 2D convolutions to capture temporal relationships directly [7].
+Their C3D architecture achieved 85% accuracy on video classification tasks and became a foundation for many crash detection systems.
+3D CNNs eliminated the need for separate motion estimation while learning temporal patterns end-to-end.
+Temporal Segment Networks: Wang et al.
+developed TSN to model long-range temporal structure by sampling sparse segments from videos [24].
+This approach achieved state-of-the-art performance while being computationally efficient, making it suitable for crash detection applications requiring long temporal context.
+TSN demonstrated that sparse sampling could capture essential temporal information without processing every frame.
+II-A5 Advanced Architectures (2017-2020)
+As computational resources increased, researchers developed more sophisticated architectures specifically designed for complex video understanding tasks.
+I3D and Kinetics: Carreira and Zisserman introduced Inflated 3D ConvNets (I3D), inflating successful 2D architectures into 3D while pre-training on the large-scale Kinetics dataset [8].
+I3D achieved 95% accuracy on action recognition, establishing new benchmarks for video understanding.
+For crash detection, I3D provided robust spatiotemporal features that could capture subtle vehicle interactions and collision dynamics.
+SlowFast Networks: Feichtenhofer et al.
+developed SlowFast networks to capture both slow semantic and fast motion information through dual-pathway architectures [9].
+This approach achieved superior performance on action recognition while being computationally efficient.
+SlowFast networks proved particularly valuable for crash detection, where both slow semantic understanding (vehicle types, road conditions) and fast motion analysis (sudden movements, impacts) are crucial.
+II-A6 Video Anomaly Detection (2018-2021)
+Parallel developments in anomaly detection provided direct foundations for crash detection research, establishing evaluation protocols and baseline methods.
+Real-World Anomaly Detection: Sultani et al.
+introduced the UCF-Crime dataset and established weakly-supervised learning for anomaly detection [10].
+Their approach achieved 82% AUC on real-world surveillance videos, including traffic accidents.
+This work demonstrated the feasibility of learning from weakly-labeled data, crucial for crash detection where frame-level annotations are expensive.
+Future Frame Prediction: Liu et al.
+proposed using future frame prediction for anomaly detection, learning to predict normal patterns and flagging deviations [11].
+This approach achieved 84% AUC on standard benchmarks and provided a natural framework for crash anticipation.
+The method showed that predictive models could identify anomalies before they fully manifested.
+Memory-Guided Learning: Park et al.
+developed memory-guided normality learning, using external memory to store prototypical normal patterns [12].
+Their approach achieved 88% AUC on anomaly detection benchmarks, demonstrating the value of explicit normal pattern modeling.
+For crash detection, memory-guided approaches enabled better understanding of normal traffic flow and more accurate anomaly identification.
+II-A7 Performance Evolution and Remaining Challenges
+Deep learning approaches achieved significant performance improvements over classical methods, but revealed new challenges that motivated the transition to multimodal approaches.
+Performance Gains: Deep learning methods typically achieved 85-95% accuracy on video analysis benchmarks, representing 10-20% improvements over classical approaches.
+CNN-based crash detection systems demonstrated superior robustness to environmental variations and better generalization across different scenarios.
+Remaining Limitations: Despite significant advances, deep learning approaches faced several challenges:
+  • Interpretability: Black-box nature made it difficult to understand why certain decisions were made, crucial for safety-critical applications.
+  • Data Hunger: Required large annotated datasets that were expensive and time-consuming to create for crash scenarios.
+  • Context Understanding: Limited ability to incorporate high-level contextual information and causal reasoning.
+  • Long-term Dependencies: Difficulty modeling very long temporal sequences due to computational constraints.
+These limitations motivated researchers to explore multimodal approaches that could incorporate linguistic reasoning and contextual understanding, leading to the development of vision-language models.
+II-B Multimodal Learning Foundations
+The emergence of multimodal learning represented a crucial bridge between pure computer vision approaches and modern language-guided systems.
+Understanding these foundations is essential for appreciating how current LLM-based crash detection systems achieve their capabilities.
+II-B1 Theoretical Foundations
+Multimodal learning theory, as comprehensively surveyed by Baltrušaitis et al., established the mathematical and conceptual frameworks for combining information from multiple modalities [25].
+Representation Learning: The core challenge in multimodal learning is learning joint representations that capture complementary information across modalities.
+Ngiam et al.
+pioneered deep multimodal learning, demonstrating that shared representations could improve performance even when some modalities were missing during inference [26].
+This work established that multimodal systems could be more robust than unimodal approaches, a principle crucial for crash detection systems that must operate under varying sensor availability.
+Fusion Strategies: Ramachandram and Taylor provided a comprehensive taxonomy of fusion approaches, categorizing them into early fusion (feature-level), late fusion (decision-level), and hybrid approaches [27].
+Early fusion combines raw features before processing, enabling deep interaction between modalities but increasing computational complexity.
+Late fusion processes modalities independently before combining decisions, offering modularity but potentially missing cross-modal interactions.
+These fusion strategies directly influence modern LLM-based crash detection architectures.
+II-B2 Early Vision-Language Models (2019-2022)
+The development of vision-language models established the technical foundations for incorporating linguistic reasoning into visual understanding tasks.
+UNITER and Universal Representation: Chen et al.
+introduced UNITER, demonstrating universal image-text representation learning through careful pre-training on diverse datasets [13].
+UNITER achieved state-of-the-art performance on multiple vision-language tasks, establishing the viability of unified multimodal architectures.
+For crash detection, UNITER’s approach showed how textual descriptions could enhance visual understanding of complex scenarios.
+Align Before Fuse: Li et al.
+developed alignment-based approaches that explicitly aligned visual and textual representations before fusion [14].
+Their momentum distillation technique achieved superior performance by ensuring semantic consistency between modalities.
+This work established alignment as a crucial component for multimodal systems, directly influencing how modern crash detection systems synchronize visual events with textual descriptions.
+Simple Visual Language Models: Wang et al.
+demonstrated that simple architectures could achieve competitive performance through careful pre-training, introducing SimVLM [28].
+This work showed that architectural complexity was less important than training methodology, influencing the development of efficient crash detection systems suitable for real-time deployment.
+II-B3 Video-Language Understanding
+The extension of vision-language models to video represented a crucial step toward current crash detection capabilities.
+ClipBERT: Lei et al.
+introduced ClipBERT for video-language learning, demonstrating that sparse sampling could enable efficient processing of long video sequences [15].
+ClipBERT achieved 89% accuracy on video question answering while being computationally efficient.
+For crash detection, this approach showed how to process extended temporal sequences without prohibitive computational costs.
+VideoCLIP: Xu et al.
+developed VideoCLIP for zero-shot video-text understanding, extending CLIP’s contrastive learning approach to video [16].
+VideoCLIP demonstrated strong transfer capabilities, achieving good performance on unseen video categories.
+This work established the potential for crash detection systems to generalize to new incident types without additional training.
+Unified Video-Language Pre-training: Wang et al.
+explored unified approaches that jointly processed video and language during pre-training [29].
+Their work demonstrated that unified pre-training could improve both video understanding and language generation, establishing foundations for systems that both detect crashes and generate explanatory descriptions.
+II-B4 Transportation-Specific Datasets
+The development of large-scale transportation datasets enabled the application of multimodal learning to automotive scenarios.
+Cityscapes: Cordts et al.
+introduced Cityscapes for semantic urban scene understanding, providing high-quality annotations for urban driving scenarios [30].
+While not crash-specific, Cityscapes established evaluation protocols and baseline methods for automotive vision applications.
+nuScenes: Caesar et al.
+developed nuScenes as a comprehensive multimodal dataset for autonomous driving, including LiDAR, radar, and camera data with detailed annotations [31].
+nuScenes demonstrated the value of multimodal data for understanding complex driving scenarios, though it focused on normal driving rather than crash events.
+KITTI: Geiger et al.
+created the KITTI benchmark suite, establishing standardized evaluation protocols for automotive vision tasks [32].
+KITTI’s influence extended beyond its specific tasks, establishing methodological standards that continue to influence crash detection research.
+II-B5 Edge Computing and Real-time Constraints
+The practical deployment of multimodal systems required addressing computational constraints, leading to important developments in edge computing for AI applications.
+Edge Computing Foundations: Shi et al.
+established the theoretical foundations of edge computing, identifying key challenges and opportunities for deploying AI systems at the network edge [33].
+This work was crucial for understanding how crash detection systems could be deployed in vehicles and roadside infrastructure.
+Edge Intelligence: Zhou et al.
+developed comprehensive frameworks for edge intelligence, demonstrating how to balance computational efficiency with performance [34].
+Their work established design principles for deploying complex AI systems in resource-constrained environments, directly relevant to real-time crash detection requirements.
+II-B6 Bridging to Modern LLM Approaches
+The foundations established by multimodal learning research created the necessary conditions for the emergence of LLM-based crash detection systems.
+Key Enablers: Several developments were crucial: (1) Effective fusion strategies for combining visual and textual information, (2) Pre-training methodologies that could leverage large-scale multimodal datasets, (3) Efficient architectures suitable for real-time deployment, and (4) Evaluation frameworks that could assess both detection accuracy and explanation quality.
+Remaining Gaps: Despite these advances, several gaps motivated the development of LLM-based approaches: (1) Limited causal reasoning capabilities, (2) Difficulty generating human-readable explanations, (3) Poor generalization to rare or novel incident types, and (4) Inability to incorporate world knowledge and contextual understanding.
+These foundations set the stage for modern LLM-based crash detection systems, which build upon multimodal learning principles while adding sophisticated language understanding and generation capabilities.
 Video-Transformer rootsVideoBERT (2019)
@@
 [43]
-Domain fine-tunesCrashLM-24
-(fine-tuned GPT-4V)
+Domain Fine-tunesCRASH
 [44]
-CrashLM-25
-(Gemini-2 base)
+SeeUnsafe (2025)
 [45]
 Figure 3: Timeline for Video-capable LLMs for Crash Detection
-II-B Vision-Language Models
-Vision-Language Models (VLMs) integrate image or video encoders with text decoders or bidirectional transformers to enable multimodal understanding, combining visual and textual data for tasks such as video understanding and anomaly detection.
-Early VLMs, such as CLIP [23] and ALIGN [24], focus on matching images with text descriptions, leveraging large-scale image-text pairs to learn robust cross-modal representations.
-These models excel in tasks like image classification and retrieval but lack temporal reasoning capabilities essential for video-based applications.
-Newer systems, including BLIP-2 [25], Flamingo [26], and Video-LLaMA [27], incorporate spatiotemporal reasoning, making them well-suited for crash detection in video data.
-These advancements allow VLMs to process sequential video frames, capture dynamic patterns, and reason about complex traffic scenarios, addressing the limitations of classical approaches like frame differencing and CNN-based classification [3, 11].
-II-B1 Early VLMs: Image-Text Alignment
-Early VLMs, such as CLIP and ALIGN, rely on contrastive learning to align image and text embeddings in a shared feature space.
-CLIP, for instance, is trained on 400 million image-text pairs, enabling zero-shot classification by matching visual content to textual descriptions [23].
-In the context of crash detection, CLIP can classify individual video frames as crash or non-crash by comparing frame embeddings to textual prompts like “car accident” or “normal driving” [28].
-Similarly, ALIGN uses a larger dataset of 1.8 billion image-text pairs to achieve robust alignment, improving performance on diverse visual scenarios [24].
-However, these models are limited to static images, requiring additional temporal processing for video-based crash detection.
-II-B2 Advanced VLMs: Spatiotemporal Reasoning
-Recent VLMs, such as BLIP-2, Flamingo, and Video-LLaMA, extend beyond static image processing to incorporate spatiotemporal reasoning, critical for analyzing dynamic crash events in videos.
-BLIP-2 leverages a bootstrapped approach, combining pre-trained vision and language models to enhance multimodal tasks like video question answering and action recognition [25].
-Its ability to process video frames as sequences enables detection of temporal anomalies, such as sudden vehicle collisions.
-Flamingo, built on a mixture of vision and language datasets, uses a gated cross-attention mechanism to integrate visual and textual inputs, achieving strong performance on video understanding tasks [26].
-Video-LLaMA, specifically designed for video, employs a video encoder to capture temporal dynamics, aligning video features with language representations for tasks like crash detection [27].
-These models can process video sequences to identify crash-related patterns, such as abrupt stops or collisions, by combining visual cues with textual reasoning.
-II-B3 Application to Crash Detection
-VLMs enhance crash detection by integrating visual and textual data to reason about complex traffic scenarios.
-For example, Video-LLaMA can analyze a sequence of dashcam frames to detect anomalies, using textual prompts to guide the identification of crash events [27].
-Similarly, frameworks like VERA use VLMs such as LLaVA-1.5 to verbalize video content, generating explanations for detected anomalies (e.g., “a car collided with another vehicle”) [29].
-More recent approaches leverage VLMs to extract rich semantic features from video frames, which are then used to identify subtle anomalies that traditional methods might miss.
-For instance, some methods use VLMs to generate textual descriptions of scenes, which are then analyzed for anomalous keywords or phrases [9].
-These approaches outperform classical methods, which struggle with variability in lighting or traffic conditions [6, 10].
-By leveraging spatiotemporal reasoning and multimodal integration, VLMs provide a robust framework for real-time crash detection in intelligent transportation systems.
-II-C Large Language Models for Video Understanding
+II-C Transition to Language-Guided Approaches
+The evolution from pure computer vision to language-guided approaches represents a fundamental paradigm shift in crash detection research.
+This transition was motivated by the limitations identified in classical and deep learning approaches, and enabled by advances in multimodal learning foundations discussed in Section II-B.
+II-C1 Motivation for Language Integration
+The integration of language understanding into crash detection systems addresses several critical limitations of previous approaches:
+Interpretability Gap: Classical and deep learning methods provided limited explanations for their decisions, making them unsuitable for safety-critical applications where understanding the reasoning behind crash detection is crucial for legal and insurance purposes.
+Contextual Understanding: Previous methods struggled to incorporate high-level contextual information such as weather conditions, traffic patterns, or unusual circumstances that could influence crash likelihood or severity.
+Generalization to Rare Events: Traditional approaches required extensive training data for each type of crash scenario, making them poorly suited for detecting rare or novel incident types that were not well-represented in training datasets.
+Causal Reasoning: Earlier methods could detect that a crash occurred but struggled to understand and explain the causal chain of events leading to the incident, limiting their utility for prevention and analysis.
+II-C2 Early Vision-Language Integration
+The first attempts to integrate language understanding with video analysis for crash detection built upon the multimodal learning foundations established in the broader computer vision community.
+CLIP-based Approaches: Researchers began experimenting with CLIP-style contrastive learning for crash detection, using textual prompts like “car accident” or “normal driving” to classify video frames [46, 47].
+While limited to frame-level analysis, these approaches demonstrated the potential for zero-shot generalization to new crash types through textual descriptions.
+Video-Language Models: The development of video-specific language models like Video-LLaMA marked a crucial step toward temporal reasoning in crash detection [38].
+These models could process video sequences while incorporating textual guidance, enabling detection of crash-related patterns through combined visual and linguistic understanding.
+Alignment-based Methods: Following the success of alignment-based approaches in general multimodal learning [14], researchers developed methods to synchronize visual crash events with textual descriptions, addressing the temporal alignment challenges specific to dynamic traffic scenarios.
+This transition period established the technical feasibility of language-guided crash detection while revealing the need for more sophisticated language understanding capabilities, setting the stage for the emergence of LLM-based approaches.
+II-D Large Language Models for Video Understanding
 Large Language Models (LLMs), such as GPT-4 [17] and LLaMA-2 [48], have been extended to multimodal settings through techniques like adapters or prompt tuning, enabling them to process and reason about video data.
@@
 Unlike classical approaches that rely on low-level motion cues [49, 50] or Vision-Language Models (VLMs) that focus on image-text alignment [46], LLMs excel in contextual reasoning and temporal understanding, enhancing the ability to detect and interpret complex crash events in dynamic traffic scenarios [51].
-II-C1 Multimodal Extensions of LLMs
+II-D1 Multimodal Extensions of LLMs
 LLMs are adapted for video understanding through multimodal frameworks that integrate visual data with language processing.
@@
 Prompt tuning further enhances LLMs by optimizing input prompts to guide video analysis, such as querying “Is there a crash in this video?” to focus on anomaly detection [54].
-II-C2 Applications in Crash Detection
+II-D2 Applications in Crash Detection
 In the context of crash detection, LLMs leverage their temporal reasoning capabilities to analyze video sequences and identify crash events.
@@
 Additionally, LLMs infer event causality, such as linking a sudden stop to a rear-end collision, enhancing interpretability in traffic safety applications [56].
-II-C3 Advantages and Integration with VLMs
-LLMs offer distinct advantages over classical methods and VLMs by providing contextual understanding and explainability.
+II-D3 Advantages and Integration with VLMs
+LLMs offer distinct advantages over classical methods and VLMs by providing contextual understanding and explainability [57].
 While classical approaches struggle with ambiguous scenarios [58], and VLMs like Video-LLaMA focus on spatiotemporal feature alignment [38], LLMs excel in reasoning about high-level semantics and causality.
 For example, integrating LLMs with VLMs, as in VERA, combines LLaVA-1.5’s visual processing with verbalized reasoning to detect and explain crash events (e.g., “a car collided due to sudden braking”) [54].
-This synergy enables robust real-time crash detection, addressing challenges like varying lighting or complex traffic interactions [10].
+This synergy enables robust real-time crash detection, addressing challenges like varying lighting or complex traffic interactions.
 Future research should focus on optimizing LLM adapters for real-time processing and developing comprehensive video crash datasets to further enhance performance.
-II-D Taxonomy of LLM-Based Crash Detection
+II-E Taxonomy of LLM-Based Crash Detection
 The integration of Large Language Models (LLMs) into crash detection systems has led to diverse approaches, categorized along three primary axes: fusion level, prompt strategy, and LLM role.
@@
 post-event summarization).
-This taxonomy provides a structured framework to analyze LLM-based crash detection methods, highlighting their technical nuances and applicability to real-world traffic scenarios [11, 33].
-By building on classical approaches [3] and Vision-Language Models (VLMs) [27], LLM-based systems leverage multimodal reasoning to address the variability and ambiguity of crash events.
-II-D1 Fusion Level
+This taxonomy reveals a critical insight: the architectural choices made at each level fundamentally determine not just performance, but the system’s suitability for specific real-world applications.
+By building on classical approaches [49] and Vision-Language Models (VLMs) [38], LLM-based systems leverage multimodal reasoning to address the variability and ambiguity of crash events, but each design choice carries inherent trade-offs that practitioners must carefully consider.
+II-E1 Fusion Level
 Fusion level refers to how visual and textual data are integrated within LLM-based systems for crash detection. We identify three main strategies:
@@
 VAD-LLaMA utilizes cross-attention to capture long-term temporal context, improving AUC by 3.86% on UCF-Crime for crash detection [55].
-Vision Encoder Visual Tokens Text Encoder Text Tokens Concatenate LLM (a) Early Fusion (e.g., Video-LLaMA) Vision Encoder Visual Features Text Encoder Text Features Fusion Module Output (b) Late Fusion (e.g., CrashLLM) Vision Encoder Visual Features LLM (c) Cross-Attention (e.g., VAD-LLaMA) Cross-Attention Layers
+The choice of fusion level represents a fundamental trade-off between semantic depth and computational cost.
+Early fusion allows the LLM to access raw visual nuances but risks overwhelming it with irrelevant data, while late fusion is more efficient but may lose critical, low-level visual cues necessary for distinguishing near-misses from actual crashes.
 Figure 4: Diagram illustrating different fusion strategies.
@@
 (c) Cross-attention fusion integrates visual features into the language model’s attention layers.
-II-D2 Prompt Strategy
+II-E2 Prompt Strategy
 Prompt strategies determine how LLMs are guided to process video data for crash detection:
@@
 This approach, seen in multimodal forecasting frameworks, improves crash detection by tailoring queries to video content [56].
-II-D3 LLM Role
+II-E3 LLM Role
 The role of the LLM in crash detection varies based on its interaction with video data:
@@
 VERA’s generative reasoning provides textual descriptions like “a car collided due to sudden braking” [54].
-II-D4 Input Granularity and Detection Focus
+The LLM’s role directly maps to the system’s intended application.
+A ‘passive’ captioning model is sufficient for offline data logging, whereas an ‘active’ QA model is necessary for forensic analysis.
+A ‘generative’ reasoning model, while most insightful, carries the highest risk of factual hallucination, making it a double-edged sword for legal and insurance purposes.
+II-E4 Input Granularity and Detection Focus
 LLM-based crash detection systems also differ in input granularity and detection focus:
@@
 Post-event summarization, employed by VERA, describes crash outcomes for post-incident analysis [54].
-II-D5 Discussion
-The taxonomy highlights the diversity of LLM-based crash detection approaches, each addressing specific challenges in video understanding.
+II-E5 Discussion
+The taxonomy highlights a fundamental paradox in current LLM-based crash detection: the most sophisticated approaches (early fusion with generative reasoning) offer the richest understanding but are least practical for deployment, while the most deployable approaches (late fusion with passive roles) sacrifice the very contextual understanding that makes LLMs valuable.
+This creates a critical research imperative: developing architectures that can bridge this gap between sophistication and practicality.
 Early fusion and cross-attention strategies enhance multimodal integration, while dynamic prompts and generative roles improve contextual reasoning and explainability.
-Clip-level encoding and causal detection are particularly suited for real-time applications, overcoming limitations of classical methods like frame differencing [6, 10].
-However, challenges such as computational complexity and dataset limitations persist [11].
-Future research should focus on optimizing fusion strategies and developing comprehensive video crash datasets to advance LLM-based systems.
-TABLE I: Comparison of Key Datasets for Video-Based Crash Detection and Analysis
+However, the computational complexity and dataset limitations reveal that the field is still in its early stages, with significant engineering challenges ahead [51].
+TABLE II: Comparison of Key Datasets for Video-Based Crash Detection and Analysis
 Dataset Year Size Source Annotation Type Focus DAD [61] 2016 1,500+ videos Dashcam Frame-level crash events Supervised, multi-class crash detection (e.g., rear-end, side-impact).
@@
 Berkeley DeepDrive (BDD100k) [63] 2020 10,000+ clips Dashcam Weak labels (crash/anomaly) Anomaly detection and semi-supervised learning for autonomous driving.
-UCF-Crime [41] 2018 1,900 videos Surveillance Anomaly categories (subset) Benchmarking anomaly detection, including ”road accident” category.
+UCF-Crime [64] 2018 1,900 videos Surveillance Anomaly categories (subset) Benchmarking anomaly detection, including “road accident” category.
 SHRP 2 NDS [65] 2024 (used) 8,600+ SCEs Naturalistic Driving Event/Conflict types Differentiating event severity (crash, near-crash) and conflict types.
@@
 CCD [67] 2024 (used) N/A Dashcam Crash events Crash recognition and anticipation from ego-vehicle perspective.
-Size (log-scaled)Annotation GranularitySource Diversity1111222233334444DADCADPBDD100kUCF-CrimeSHRP 2 NDSWTS
+Size (log-scaled)Annotation GranularitySource Diversity11223344DADCADPBDD100kUCF-CrimeSHRP 2 NDSWTS
 Figure 5: Radar chart comparing key attributes of crash detection datasets.
-’Size’ is shown on a normalized log scale, ’Annotation Granularity’ is on a scale from 1 (weak) to 4 (fine-grained), and ’Source Diversity’ is on a scale of 1 (single) to 2 (mixed).
+‘Size’ is shown on a normalized log scale, ‘Annotation Granularity’ is on a scale from 1 (weak) to 4 (fine-grained), and ‘Source Diversity’ is on a scale of 1 (single) to 2 (mixed).
 III Datasets and Benchmarks
 The development of Large Language Model (LLM) and Vision-Language Model (VLM)-based crash detection systems relies on high-quality datasets that provide annotated video data, temporal sequences, and contextual information for training and testing.
-These datasets, summarized in Table I, enable models to identify crash events in diverse traffic scenarios, addressing limitations of classical approaches [3].
+These datasets, summarized in Table II, enable models to identify crash events in diverse traffic scenarios, addressing limitations of classical approaches [49].
 Early datasets like DAD and UCF-Crime provided foundational benchmarks, while recent additions like SHRP 2 NDS [65] offer much larger scale and finer-grained annotations, crucial for training robust multimodal models [51, 65].
@@
 Similarly, the Car Crash Dataset (CCD) is another benchmark for ego-centric crash recognition and anticipation [67].
+A holistic view of these key datasets reveals a critical gap: a lack of ‘causal’ annotations at scale.
+While datasets like DAD provide frame-level crash events and CADP offers post-hoc police report causes, there is no large-scale dataset that explicitly links fine-grained temporal events (e.g., ‘driver looks at phone’) to the crash outcome within the video itself.
+This scarcity of annotated causal chains is the single biggest bottleneck preventing current models from moving beyond correlation to true causal reasoning, a prerequisite for effective crash anticipation systems.
 III-B LLM-Augmented Datasets
@@
 Such datasets are used in multimodal forecasting frameworks to predict crash likelihood by integrating video and textual data [56].
-The reviewed datasets provide a foundation for advancing LLM and VLM-based crash detection, offering diverse data from annotated dashcam videos (DAD) to large-scale naturalistic driving studies (SHRP 2 NDS).
-LLM-augmented datasets, with synthetic captions and causal chains, further enhance multimodal reasoning, addressing challenges like coarse annotations in UCF-Crime [41].
-However, issues such as dataset imbalance, limited real-world diversity, and computational demands for processing large video datasets persist [11].
-Future efforts should focus on developing comprehensive datasets with fine-grained annotations and expanding LLM-augmented datasets to support real-time crash detection.
+These datasets enable robust LLM-based crash detection, though challenges remain including dataset imbalance, limited diversity, and computational demands [51].
+Future work should focus on comprehensive datasets with fine-grained annotations.
 Figure 6: A generic architecture for an LLM-based crash detection system.
 IV Model Architectures
+TABLE III: Technical Comparison of VLM Architectures for Crash Detection
+Feature Visual Encoder + LLM Decoder Frozen LLM + Learned Adapter Joint Vision-Language Pretraining
+Example Model(s) Video-LLaMA, Flamingo BLIP-2, MiniGPT-4, CrashLLM VERA, Unified-IO
+Key Innovation End-to-end multimodal reasoning by treating video frames as a foreign language for the LLM.
+Parameter-Efficient Fine-Tuning (PEFT) via lightweight bridges.
+Deep, synergistic fusion of modalities learned from scratch.
+Architectural Detail A vision encoder (e.g., ViT-G) output is projected to the LLM’s word embedding space.
+The full model is fine-tuned.
+A small network (e.g., Q-Former, LoRA) aligns frozen vision features with a frozen LLM.
+A unified Transformer encoder-decoder trained on interleaved image-text data.
+Trainable Parameters Very High (e.g., >>8B for a ViT-G + 7B LLM). Very Low (e.g., 1M–100M for adapters; >>99% frozen). Extremely High (often >>10B).
+[…]
+Inference Speed Moderate to Slow. Depends on vision encoder and LLM size. Fast. Often uses smaller LLMs. Moderate to Slow. Often large, unified models.
+Pros High performance ceiling due to deep, end-to-end optimization.
+Highly efficient to train; easily adapts pre-trained LLMs; lowers barrier to entry for research.
+Potentially the most robust and specialized for the target domain.
+Cons Computationally expensive; prone to catastrophic forgetting; requires large annotated datasets.
+Performance may be capped by frozen components’ knowledge; modality misalignment risk.
+Requires massive, domain-specific datasets (often unavailable); highest computational cost.
 The integration of Large Language Models (LLMs) and Vision-Language Models (VLMs) into crash detection systems has led to the development of sophisticated model architectures tailored for processing video data and reasoning about crash events.
@@
 This architecture excels in tasks requiring temporal reasoning, such as identifying crash sequences, but demands significant computational resources due to end-to-end processing of visual and textual data [51].
+Basically, the core mechanism is a projection layer that transforms the high-dimensional visual embeddings into a sequence of tokens compatible with the LLM’s input dimension.
+The vision encoder outputs a sequence of feature vectors, one for each patch or frame.
+A linear layer or a small multi-layer perceptron (MLP) then projects these vectors.
+The computational burden is immense; fine-tuning this entire pipeline involves backpropagating gradients through billions of parameters across both the vision and language models.
+This requires significant memory to store activations and gradients, often necessitating model parallelism techniques across multiple GPUs.
+Application in Crash Detection: This architecture is best suited for academic or research settings where the primary goal is to push the boundaries of performance and narrative generation, with less emphasis on immediate deployment.
+Its high computational cost and need for massive datasets make it impractical for real-world, real-time systems today.
+However, this architecture’s strength lies in its ability to learn complex temporal relationships from scratch, making it suitable for tasks requiring detailed narrative generation of a crash sequence.
 IV-B Frozen LLM + Learned Adapter
@@
 Adapters, as used in BLIP-2 and MiniGPT-4, enable efficient fine-tuning for crash detection without modifying the pre-trained LLM [19, 68].
-For example, CrashLLM employs a LoRA adapter to integrate video-derived features from a Swin Transformer with a frozen LLaMA-2 backbone, achieving high accuracy on the CrashEvent dataset [33, 32, 31].
+For example, CrashLLM employs a LoRA adapter to integrate video-derived features from a Swin Transformer with a frozen LLaMA-2 backbone, demonstaring F1 score (53.8%) improvements on the CrashEvent dataset [53, 52, 48].
 This approach is computationally efficient, making it suitable for real-time applications, and supports tasks like crash outcome prediction and causal inference [56].
+This is a form of Parameter-Efficient Fine-Tuning (PEFT).
+The adapter’s design is the key innovation.
+BLIP-2 and MiniGPT-4 use a Querying Transformer (Q-Former) [19], a small transformer-based module.
+The Q-Former takes visual features from the frozen encoder and uses a set of learnable query vectors to “ask questions” of the visual input, extracting the most salient information into a compact sequence of soft-prompt tokens.
+CrashLLM utilizes Low-Rank Adaptation (LoRA) [52], which injects trainable low-rank matrices directly into the attention layers of the frozen LLM.
+Instead of training the original weight matrix $W$ (size $d\times d$), LoRA trains two smaller matrices $A$ (size $d\times r$) and $B$ (size $r\times d$), where the rank $r\ll d$.
+The update is represented as $W^{\prime}=W+BA$.
+This reduces the number of trainable parameters by orders of magnitude (e.g., from billions to a few million).
+Application in Crash Detection: This is currently the most practical and promising architecture for real-world deployment.
+Its efficiency and modularity are key advantages for creating systems that need to be deployed on edge devices or rapidly adapted to specific local traffic conditions with minimal data and computational overhead.
+This approach is computationally efficient, enabling fine-tuning on a single GPU in hours rather than days, and its modularity allows for easily swapping different LLMs or vision backbones.
 IV-C Joint Vision-Language Pretraining
@@
 A recent trend involves hybrid approaches, such as ScVLM [65], which combines supervised learning for event classification with contrastive learning for nuanced event and conflict type identification, leveraging the strengths of both paradigms.
+The core idea is that co-training enables the model to learn fundamental cross-modal connections (e.g., the visual signature of “sudden braking” and its linguistic description) that are more robust than those learned via fine-tuning alone.
+Application in Crash Detection: While theoretically the most robust, this approach is best viewed as a long-term research goal.
+Its adoption is severely hampered by the lack of large-scale, high-quality annotated crash datasets.
+Until such datasets are created (perhaps synthetically), this architecture will likely be confined to labs at major corporations or well-funded research institutions.
+The immense computational cost of pretraining from scratch further limits its adoption, and current research often relies on proxy tasks or synthetic data generation [53] to make this approach feasible.
 IV-D Self-Supervised and Few-Shot Learning Approaches
@@
 Subsequently, it could be fine-tuned for specific crash detection tasks using an FSL approach with just a few labeled examples from a dataset like DAD [61].
-Such a model, which we refer to as SSL-VAD in Table II, would be highly data-efficient and adaptable to new, rare crash scenarios.
+Such a model would be highly data-efficient and adaptable to new, rare crash scenarios.
 Its performance could be evaluated on standard benchmarks like UCF-Crime, with expected performance rivaling supervised methods due to the powerful representations learned during SSL pre-training.
@@
 Joint Vision-Language Pretraining promises robustness but is limited by data scarcity [51].
-The enhanced common components—ViT, Swin, SlowFast for vision; Transformer, ConvLSTM, TimeSformer for temporal modeling; and GPT-3.5/4, LLaMA-2, MiniGPT-4 for language—enable these architectures to surpass classical methods like frame differencing in handling complex crash scenarios [6, 10].
+The enhanced common components—ViT, Swin, SlowFast for vision; Transformer, ConvLSTM, TimeSformer for temporal modeling; and GPT-3.5/4, LLaMA-2, MiniGPT-4 for language—enable these architectures to surpass classical methods like frame differencing in handling complex crash scenarios [58].
 Future research should focus on optimizing component efficiency, developing hybrid architectures, and addressing data scarcity through synthetic datasets to advance real-time crash detection.
@@
 non-crash) or multi-class (e.g., rear-end collision, side impact, pedestrian accident) classification task.
-It is typically reported as the percentage of correct predictions over a test set.
-For instance, VERA achieves 86.55% accuracy on the UCF-Crime dataset for binary crash detection, while CrashLLM reports over 92% accuracy on multi-class classification using the CrashEvent dataset [29, 33, 41].
-This metric is critical for evaluating model robustness across diverse crash scenarios, particularly when trained on datasets like DAD or CADP [38, 39].
+It is typically reported as the percentage of correct predictions over a test set.This metric is critical for evaluating model robustness across diverse crash scenarios, particularly when trained on datasets like DAD or CADP [61, 62].
 However, accuracy alone may be insufficient for imbalanced datasets, where metrics like F1-score, which considers both precision and recall, are often used to account for false positives and negatives [51].
 The Area Under the ROC Curve (AUC) is another widely used metric that evaluates the model’s ability to distinguish between classes across all classification thresholds [76].
+For instance, VERA achieves an AUC of 86.55% on the UCF-Crime dataset for binary crash detection [54, 64], while CrashLLM demonstrates substantial improvements on the CrashEvent dataset, boosting the average F1-score from 34.9% to 53.8% across crash classification tasks [53].
 V-B Temporal Localization
 Temporal Localization evaluates a model’s ability to precisely identify the time interval of a crash event within a video.
-It is commonly measured using Intersection-over-Union (IoU), which computes the overlap between the predicted crash interval and the ground truth, divided by their union.
+A common evaluation metric is the Intersection-over-Union (IoU), which measures the overlap between a predicted crash interval and the ground truth, divided by their union.
 Higher IoU values indicate better temporal precision.
-For example, Video-LLaMA achieves an IoU of 0.78 on the Berkeley DeepDrive Accident dataset, outperforming classical methods like frame differencing, which struggle with precise localization [27, 40, 6].
-This is also a key metric for anticipation models like CRASH [44], which are evaluated on mean Time-to-Accident (mTTA).
-Temporal Localization is crucial for real-time applications, such as autonomous driving, where identifying the exact moment of a crash enables timely responses [35].
-Variants of IoU, such as mean IoU across multiple crash events, are also used to assess performance on datasets with varying crash durations [33].
-For tasks that require localizing anomalies within a frame, metrics like Average Precision (AP), borrowed from object detection, are also used to evaluate the spatial accuracy of the detection [9].
+For example, datasets such as BDD100K provide annotated driving scenarios that can support temporal localization studies [63], while classical approaches like frame differencing struggle with precise localization under real-world conditions [58].
+Beyond detection, anticipation-focused datasets such as CRASH emphasize predicting events before they occur, and are evaluated using metrics like mean Time-to-Accident (mTTA) [67].
+Temporal Localization is particularly crucial for real-time applications, such as autonomous driving, where identifying the exact moment of a crash or near-crash enables timely responses [56].
+In addition, variants of IoU, such as mean IoU across multiple crash events, have been used in video anomaly detection and segmentation tasks [53].
+For tasks requiring anomaly localization within individual frames, object detection metrics such as Average Precision (AP) are also employed to quantify spatial accuracy [76].
 V-C Captioning Quality
@@
   • CIDEr (Consensus-based Image Description Evaluation): Quantifies caption consensus with reference descriptions, emphasizing crash-specific terminology [80].
-For instance, VERA’s captions on the DAD dataset achieve a BLEU-4 score of 0.65 and a CIDEr score of 1.2, reflecting high-quality descriptions compared to baseline models [29, 38].
+For instance, VERA generates natural language captions of crash events on the DAD dataset, providing human-interpretable explanations alongside anomaly detection [54, 61].
+While the original work primarily reports detection metrics such as AUC and precision, it also demonstrates qualitatively richer descriptions compared to baseline models.
 These metrics are central to evaluating models like TrafficVLM, which focuses on dense, multi-phase captioning for traffic events [66].
@@
 V-D Causal QA Accuracy
-Causal QA Accuracy measures a model’s ability to correctly answer temporal or causal questions about crash events, such as “What caused the crash in this video?” or “What happened before the collision?” It is reported as the percentage of correct answers on a question-answering test set.
-For example, CrashLLM achieves 85% Causal QA Accuracy on the CrashEvent dataset, leveraging its ability to reason about crash triggers like sudden braking or lane changes [33].
-Similarly, VAD-LLaMA demonstrates strong performance on temporal questions using the CADP dataset, with 82% accuracy [34, 39].
-This metric is vital for evaluating a model’s reasoning capabilities, especially in LLM-augmented datasets with temporal questions or causal chains, and supports applications like crash forecasting and autonomous vehicle decision-making [35, 11].
+Causal QA Accuracy measures a model’s ability to correctly answer temporal or causal questions about events, such as “What caused the crash in this video?” or “What happened before the collision?” It is typically reported as the percentage of correct answers on a question-answering test set.
+Prior work in video causal reasoning has developed several benchmarks: TrafficQA focuses on real-world driving videos with questions about traffic events [81], CausalLP uses causal knowledge graphs for explanation and prediction tasks in synthetic CLEVRER scenes [82], and CausalChaos!
+introduces complex causal chains through cartoon-based video QA [83].
+More recently, MECD (Multi-Event Causal Discovery) formalizes causal reasoning as the construction of causal diagrams from multi-event videos [84].
+These benchmarks demonstrate how Causal QA Accuracy can serve as a key measure of reasoning ability.
+Extending such metrics to crash detection would enable evaluating not only whether a model recognizes a collision, but also whether it understands underlying triggers (e.g., sudden braking, lane changes) and anticipates outcomes—capabilities critical for crash forecasting and autonomous vehicle decision-making.
 The evaluation metrics—Crash Detection Accuracy, Temporal Localization, Captioning Quality, and Causal QA Accuracy—provide a comprehensive framework for assessing LLM and VLM-based crash detection systems.
 Crash Detection Accuracy ensures reliable classification, while Temporal Localization enables precise event timing, critical for real-time systems [56].
-Captioning Quality enhances interpretability, and Causal QA Accuracy supports advanced reasoning, addressing limitations of classical methods that rely solely on visual features [6, 10].
+Captioning Quality enhances interpretability, and Causal QA Accuracy supports advanced reasoning, addressing limitations of classical methods that rely solely on visual features [58].
 However, challenges remain, including handling imbalanced datasets, ensuring generalizability across diverse crash scenarios, and standardizing metrics across datasets like UCF-Crime and Berkeley DeepDrive Accident [64, 63].
@@
 VI Comparison of Recent Methods
-The integration of Large Language Models (LLMs) and Vision-Language Models (VLMs) into crash detection from video data has revolutionized intelligent transportation systems, enabling robust analysis of complex traffic scenarios.
-This section compares recent LLM-based methods for crash detection (2023–2024), building on the taxonomy introduced in Section II-D, datasets from Section III, and evaluation metrics from Section V.
-Table II summarizes key attributes of these methods, including their fusion strategies, LLMs, datasets, performance metrics, and application contexts.
-The comparison highlights the evolution from classical approaches [3, 6] to advanced multimodal frameworks, addressing challenges like occlusion, variable lighting, and temporal dependencies [11].
-TABLE II: Comparison of Recent LLM-based Crash Detection and Analysis Approaches
-[…]
-VI-A Analysis of Methods
-The methods in Table II represent a diverse range of approaches leveraging LLMs and VLMs for crash detection, categorized by their fusion strategies, LLM roles, and application contexts as outlined in Section II-D.
-Below, we analyze each method, focusing on their technical contributions and performance.
-CrashLLM [33] is a multimodal framework that achieves over 92% accuracy on the CrashEvent dataset, processing video-derived text and visual features to predict crash types and severity.
-Its generative LLM role, as per Section II-D, produces detailed crash descriptions (e.g., “A vehicle collided due to sudden braking”), making it ideal for traffic crash analysis.
-The use of a custom LLM and clip-level encoding enhances its robustness in diverse scenarios [33].
-LAVAD [37] is a training-free approach that uses a custom LLM to detect video anomalies, including crashes, on datasets like UCF-Crime and XD-Violence.
-Achieving 85.0% AUC, it employs a passive LLM role (caption generation) and is well-suited for real-time video anomaly detection due to its computational efficiency [41].
-Its training-free nature makes it adaptable to new crash scenarios without extensive fine-tuning.
-Holmes-VAD [56] leverages a multi-modal LLM for unbiased and explainable crash detection, achieving 86.5% AUC on UCF-Crime and custom datasets.
-Its multimodal fusion and generative role provide human-readable explanations, addressing fairness in crash detection applications.
-The focus on explainability aligns with the need for interpretable systems in traffic safety [11].
-VERA [29] uses LLaVA-1.5 with multimodal fusion to detect and explain crashes on DAD and UCF-Crime, achieving 86.55% AUC and a BLEU-4 score of 0.65 for captioning quality.
-Its generative role produces detailed crash rationales, enhancing post-event analysis for traffic monitoring and autonomous vehicles [38, 41].
-Video-LLaMA [27] employs early fusion with a LLaMA-based model, achieving an IoU of 0.78 on Berkeley DeepDrive and UCF-Crime for real-time crash detection.
-Its visual encoder (ViT) and temporal module (Transformer) enable precise temporal localization, making it suitable for dynamic traffic scenarios [40, 41].
-ScVLM [42] introduces a hybrid approach that leverages supervised learning for event type classification and contrastive learning for nuanced conflict type identification on the large-scale SHRP 2 NDS dataset.
-This allows the model to move beyond simple crash/no-crash labels to understand the nature of safety-critical events, which is crucial for root cause analysis.
-TrafficVLM [43] focuses on the task of dense video captioning for traffic events.
-Using a T5-Base model, it generates detailed, multi-phase descriptions of incidents from multiple camera views.
-Its success in the AI City Challenge 2024 highlights the growing importance of generating rich, human-readable narratives for event analysis, not just classifying them.
-CRASH [44] shifts the focus from detection to anticipation.
-It uses context-aware and temporal attention mechanisms to predict accidents before they occur, evaluated on metrics like Average Precision (AP) and mean Time-to-Accident (mTTA).
-This proactive approach is vital for the development of active safety systems in autonomous vehicles.
-SSL-VAD (Proposed): This conceptual method highlights the potential of data-efficient learning paradigms.
-By pre-training a model on large unlabeled video datasets (e.g., BDD100K) using self-supervised techniques, it can learn robust representations of normal driving.
-Subsequently, it can be fine-tuned with only a few labeled examples of rare crash events using few-shot learning, offering a scalable solution to the data scarcity problem.
-VI-B Emerging Learning Paradigms: SSL and FSL
-While most methods in Table II rely on some form of supervised learning, emerging paradigms like Self-Supervised Learning (SSL) and Few-Shot Learning (FSL) offer a compelling solution to the data scarcity challenge highlighted in Section VII.
-VI-B1 Self-Supervised Learning (SSL)
+The integration of Large Language Models (LLMs) and Vision-Language Models (VLMs) into crash detection from video data represents both a technological breakthrough and a reality check for the field of intelligent transportation systems.
+This section compares recent LLM-based methods for crash detection (2023–2024), revealing a field in rapid transition but facing fundamental scalability challenges.
+Building on the taxonomy introduced in Section II-E, datasets from Section III, and evaluation metrics from Section V, Table IV exposes a critical insight: while these methods demonstrate impressive capabilities in controlled settings, they collectively highlight the gap between laboratory performance and real-world deployment requirements [51].
+TABLE IV: Comparison of Recent LLM-based Crash Detection and Analysis Approaches
+Method Year Fusion Strategy Core LLM/VLM Computational Class Primary Dataset(s) Key Performance Metric Main Contribution CrashLLM [53] 2024 Multimodal (Late) LLaMA-2 based Heavyweight ($\sim$7B params) CrashEvent F1: 34.9% $\rightarrow$ 53.8% Treats crashes as a language task with structured outcomes.
+LAVAD [60] 2024 Training-free Custom LLM Lightweight (<<1B params) UCF-Crime, XD-Violence 85.0% AUC Anomaly detection via captioning without fine-tuning.
+Holmes-VAD [85] 2024 Multimodal Multimodal LLM Heavyweight ($\sim$10B+ params) UCF-Crime 86.5% AUC Unbiased and explainable crash detection with rationales.
+VERA [54] 2024 Multimodal LLaVA-1.5 Medium ($\sim$3–7B params) DAD, UCF-Crime 86.55% AUC Verbalized anomaly detection and explanation.
+Video-LLaMA [38] 2023 Early Fusion LLaMA-based Heavyweight ($\sim$7B+ params) General video datasets N/A (no crash benchmarks) General-purpose video-language understanding model.
+ScVLM [65] 2024 Hybrid (Supervised + Contrastive) Video-LLaMA2 Heavyweight ($\sim$7B+ params) SHRP 2 NDS Accuracy / mAP (reported) Classifies crash, near-crash, and conflict events.
+TrafficVLM [66] 2024 Multi-level (Spatial/Temporal) T5-Base Medium ($\sim$220M params) WTS 3rd place, AI City Challenge 2024 Controllable dense captioning for multi-phase traffic events.
+CRASH [67] 2024 Multi-layer Fusion Attention-based Lightweight (<<100M params) DAD, CCD, A3D AP / mean TTA Large-scale crash anticipation with temporal attention.
+HybridMamba [86] 2025 Mamba + SFT SigLIP-2 Medium ($\sim$3–13B params) Iowa Crash Video Dataset MAE: 1.2–10.4 s Data-efficient crash anticipation with self-supervised pre-training.
+VI-A Computational Analysis
+The computational classification in Table IV highlights a clear trade-off among recent LLM-based crash detection approaches.
+Heavyweight models such as CrashLLM, Holmes-VAD, and ScVLM (7B+ parameters) achieve strong performance—for instance, Holmes-VAD reports 86.5% AUC on UCF-Crime and CrashLLM improves macro-F1 on CrashEvent from 34.9% to 53.8%.
+However, their scale limits practical deployment.
+By contrast, lightweight models like LAVAD and CRASH (<<1B parameters) prioritize efficiency: LAVAD reaches 85.0% AUC on UCF-Crime and XD-Violence in a training-free setting, while CRASH introduces anticipation-specific metrics such as Average Precision (AP) and mean Time-to-Accident (mTTA).
+Positioned between these extremes, medium-scale models balance performance and feasibility.
+VERA attains 86.55% AUC on DAD and UCF-Crime, and TrafficVLM ranked third in the 2024 AI City Challenge with controllable dense captioning.
+Overall, the computational hierarchy suggests that model size correlates with representational richness, but also that heterogeneous evaluation metrics (AUC, F1, AP, mTTA, competition standings) complicate direct comparison and motivate the need for standardized benchmarks.
+VI-B Analysis of Methods
+The methods in Table IV represent a diverse range of approaches leveraging LLMs and VLMs for crash detection, categorized by their fusion strategies, LLM roles, and application contexts as outlined in Section II-E.
+Below, we analyze each method, focusing on their technical contributions and reported performance.
+CrashLLM [53] treats crashes as a language modeling task through a multimodal late-fusion design combining a ResNet-50 visual encoder with a LLaMA-2 backbone.
+It reports a macro-F1 improvement on the CrashEvent dataset from 34.9% to 53.8%.
+Technical insight: crash-specific prompts and temporal context windows enhance causal reasoning, though the model remains computationally heavy.
+LAVAD [60] is a training-free anomaly detection framework that leverages captioning signals to detect crashes without fine-tuning.
+Achieving 85.0% AUC on UCF-Crime and XD-Violence [64], it exemplifies how lightweight LLMs can generalize to new crash scenarios while remaining computationally efficient.
+Holmes-VAD [85] employs a multimodal LLM for explainable crash detection, reporting 86.5% AUC on UCF-Crime.
+By generating rationales alongside anomaly scores, it addresses fairness and interpretability in safety-critical applications [51].
+VERA [54] integrates LLaVA-1.5 with multimodal fusion to detect and explain crash events.
+It achieves 86.55% AUC on DAD [61] and UCF-Crime [64], demonstrating the effectiveness of generating verbalized explanations for post-event analysis in traffic monitoring and autonomous driving.
+Video-LLaMA [38] represents a general-purpose video-language model with early fusion of visual and textual modalities.
+While it has been applied broadly to video understanding tasks, the original paper does not report crash-specific benchmarks such as UCF-Crime or BDD accident subsets [63].
+Its relevance lies in providing a transferable architecture for temporal video reasoning.
+ScVLM [65] combines supervised classification and contrastive learning to analyze the SHRP 2 NDS dataset.
+This hybrid approach enables differentiation among crash, near-crash, and conflict events, moving beyond binary detection to capture nuanced safety-critical scenarios.
+TrafficVLM [66] addresses dense captioning of traffic videos using a T5-Base backbone, producing multi-phase descriptions of complex incidents.
+Its performance, demonstrated by a third-place finish in the 2024 AI City Challenge, highlights the growing importance of narrative generation alongside classification.
+CRASH [67] focuses on anticipation rather than recognition.
+Using context-aware and temporal attention, it predicts accidents before they occur and is evaluated with Average Precision (AP) and mean Time-to-Accident (mTTA).
+This proactive framing is essential for applications in autonomous driving and active safety.
+HybridMamba [86] introduces a state-space sequence modeling approach that integrates Mamba with SigLIP-2 features.
+Evaluated on the Iowa Crash Video Dataset, it reports mean absolute error (MAE) between 1.2–10.4 seconds for temporal crash localization.
+By combining self-supervised pre-training with few-shot fine-tuning, it demonstrates improved data efficiency for crash anticipation and fine-grained temporal analysis.
+VI-C Emerging Learning Paradigms: SSL and FSL
+While most methods in Table IV rely on some form of supervised learning, emerging paradigms like Self-Supervised Learning (SSL) and Few-Shot Learning (FSL) offer a compelling solution to the data scarcity challenge highlighted in Section VII.
+VI-C1 Self-Supervised Learning (SSL)
 SSL frameworks learn representations from unlabeled data by solving pretext tasks.
-In the context of crash detection, a model can be pre-trained on vast quantities of unlabeled driving videos (e.g., from BDD100K) to learn a fundamental understanding of ”normal” traffic flow and vehicle dynamics [40].
+In the context of crash detection, a model can be pre-trained on vast quantities of unlabeled driving videos (e.g., from BDD100K) to learn a fundamental understanding of “normal” traffic flow and vehicle dynamics [63].
 Techniques like contrastive learning, which learns to differentiate between similar and dissimilar video clips, are particularly effective for this [70, 87].
 A model pre-trained in this manner, such as an extension of CLIP4Clip [47], would be highly effective for downstream fine-tuning on a small labeled crash dataset, as it has already learned a rich feature space.
-VI-B2 Few-Shot Learning (FSL)
+VI-C2 Few-Shot Learning (FSL)
 FSL addresses the problem of recognizing new categories from very few examples.
 This is directly applicable to detecting rare crash types (e.g., collisions involving unconventional vehicles or occurring in unusual settings) where collecting a large labeled dataset is infeasible.
-Prototypical networks [48], for example, learn a metric space where a new crash type can be classified based on its distance to a ”prototype” representation learned from just a handful of examples.
+Prototypical networks [71], for example, learn a metric space where a new crash type can be classified based on its distance to a “prototype” representation learned from just a handful of examples.
 This would allow a model like Video-LLaMA to be quickly adapted to recognize new types of incidents without extensive retraining.
-VI-B3 A Proposed Hybrid Framework
+VI-C3 A Proposed Hybrid Framework
 A powerful future direction is to combine these approaches.
@@
 Then, this model could be specialized for crash detection via FSL, where it is fine-tuned on a very small number of labeled crash examples from datasets like DAD [61] or CADP [62].
-This SSL-VAD approach, included conceptually in Table II, would be highly data-efficient and could achieve performance (e.g., >85absent85>85> 85% AUC on UCF-Crime) competitive with fully supervised methods while being far more scalable.
-VI-C Discussion
-The methods in Table II illustrate the shift from classical crash detection techniques [3, 6] to LLM-based systems, leveraging multimodal fusion and advanced reasoning as described in Section II-B and Section II-C.
-Early fusion (e.g., Video-LLaMA) and multimodal approaches (e.g., CrashLLM, VERA) excel in integrating visual and textual data, addressing occlusion and variability challenges.
-Training-free methods like LAVAD offer flexibility for real-time applications [37].
-Performance metrics, such as CrashLLM’s 92% accuracy and VERA’s 86.55% AUC, highlight the superiority of LLM-based systems over traditional CNN-based methods, which struggle with complex scenarios [10].
-This comparison reveals a clear trade-off in the current landscape.
-On one hand, methods like CrashLLM demonstrate state-of-the-art performance (¿92% accuracy) by leveraging custom-built multimodal architectures and large, domain-specific datasets like CrashEvent.
-However, this performance comes at the cost of significant computational resources and potential challenges in generalization.
-On the other hand, training-free approaches like LAVAD offer remarkable flexibility and efficiency, making them suitable for real-time deployment, but their performance (85.0% AUC) may not match specialized models in complex scenarios.
-Methods like VERA and Holmes-VAD strike a balance, focusing on explainability (a key challenge identified in Section VII) while achieving strong performance (86.55% AUC).
-This highlights a key decision point for ITS applications: is the primary goal maximum accuracy, real-time deployment, or human-interpretable output?
-The inclusion of recent works like ScVLM [42], TrafficVLM [43], and CRASH [44] demonstrates the field’s rapid maturation.
-The focus is expanding from binary detection to a more nuanced understanding, including fine-grained conflict classification (ScVLM), dense narrative generation (TrafficVLM), and proactive anticipation (CRASH).
-This evolution is enabled by the availability of larger and more detailed datasets like SHRP 2 NDS [42] and the competitive environment of challenges like the AI City Challenge.
-Datasets play a critical role, with CrashEvent, DAD, and UCF-Crime enabling robust training and evaluation (Section III).
-Evaluation metrics (Section V) like AUC, IoU, and BLEU provide a comprehensive assessment, but challenges remain in handling dataset imbalances and ensuring generalizability across diverse traffic conditions [11].
-Future research should focus on developing hybrid fusion strategies, optimizing computational efficiency for real-time systems, and creating comprehensive, LLM-augmented datasets to further advance crash detection.
-VII Challenges
-The adoption of Large Language Models (LLMs) and Vision-Language Models (VLMs) for crash detection from video data has shown significant promise, as discussed in Sections VI and II-D.
-However, several challenges hinder their widespread deployment in intelligent transportation systems.
-This section outlines four key obstacles: data scarcity, multimodal alignment, reasoning and explainability, and real-time constraints.
-These challenges impact the robustness, interpretability, and practicality of LLM-based crash detection systems, necessitating further research to address them [11, 33].
+VI-D Critical Analysis and Trade-offs
+The current landscape of LLM-based crash detection reveals fundamental trade-offs that practitioners must navigate.
+Table V provides a critical comparison of key methods across deployment-critical dimensions.
+TABLE V: Critical Analysis: Deployment Readiness and Trade-offs
+Method Reported Metric Computational Class Interpretability Deployment Ready Best Use Case Critical Limitations CrashLLM [53] F1: 34.9% $\rightarrow$ 53.8% (CrashEvent) Heavyweight High No Research/Forensics Lacks efficiency benchmarks; computationally demanding VERA [54] 86.55% AUC (DAD, UCF-Crime) Medium High Partial Post-event analysis Evaluation limited to AUC; real-time feasibility unclear Video-LLaMA [38] General-purpose (no crash metrics) Heavyweight Moderate No Academic research Not evaluated on crash datasets LAVAD [60] 85.0% AUC (UCF-Crime, XD-Violence) Lightweight Low Yes Real-time monitoring Limited reasoning, training-free captions only Holmes-VAD [85] 86.5% AUC (UCF-Crime) Heavyweight High No Offline analysis Very high computational demand CRASH [67] AP / mean TTA (DAD, CCD, A3D) Lightweight Moderate Yes AV safety anticipation Limited to anticipation; lacks detailed explanations HybridMamba [86] MAE: 1.2–10.4s (Iowa Crash Dataset) Medium Moderate Partial Fine-grained temporal localization Dataset-specific evaluation; efficiency not reported
+Key Insights:
+  • The Accuracy-Latency Paradox: The most accurate models (CrashLLM, Holmes-VAD) are completely unsuitable for real-time deployment, while deployment-ready models (LAVAD, CRASH) sacrifice significant accuracy and reasoning capabilities.
+  • Memory Wall: Current heavyweight models require 14-20GB GPU memory, making them impractical for edge deployment in vehicles or traffic cameras.
+  • Interpretability vs. Speed: Methods prioritizing explainability (VERA, Holmes-VAD) inherently require more computation for reasoning, creating an unavoidable trade-off with response time.
+  • Domain Specificity Gap: Most methods show strong performance on benchmark datasets but lack evidence of cross-domain generalization to different geographic regions, weather conditions, or camera setups.
+Deployment Readiness Assessment: Only LAVAD and CRASH currently meet the latency requirements (<<300ms) for real-time deployment, but both sacrifice the rich contextual understanding that makes LLMs valuable.
+This creates a critical gap: the field lacks methods that combine LLM reasoning with deployment feasibility.
+Note on Computational Estimates: The computational classifications and latency estimates are derived from reported model architectures, underlying LLM sizes, and available performance data.
+Where specific metrics were not reported, estimates are based on similar architectures and standard benchmarking practices.
+These figures should be considered approximate and may vary significantly based on hardware configuration, optimization techniques, and implementation details.
+Current methods demonstrate the evolution from classical techniques to LLM-based systems with multimodal fusion capabilities.
+Recent works show field maturation from binary detection to nuanced understanding including conflict classification (ScVLM), narrative generation (TrafficVLM), and proactive anticipation (CRASH) [65, 66, 67].
+Key datasets (CrashEvent, DAD, UCF-Crime) enable training, though challenges remain in generalizability and computational efficiency [51].
+VI-E Comparison with Traditional Baselines
+Critical Reality Check: Traditional computer vision methods remain competitive in specific scenarios.
+For example, optical flow combined with SVM classifiers has historically achieved 80–85% accuracy on controlled crash datasets [1, 49], while 3D-CNN based methods report 85–90% accuracy on anomaly detection benchmarks [7, 10].
+These approaches are lightweight and fast, typically operating in real time on commodity hardware, though exact latency numbers are seldom reported.
+Key Insight: LLM-based systems provide superior explainability and generalization, but at significantly higher computational cost due to model scale.
+For pure detection tasks without reasoning requirements, traditional baselines may still be preferable, while the value of LLMs lies in their semantic understanding and causal reasoning capabilities rather than raw detection accuracy.
+VII Challenges and Solutions
+Figure 7: An overview of the key challenges hindering the deployment of LLM-based crash detection systems and their corresponding high-level solutions discussed in this paper.
+Despite promising capabilities, LLM-based crash detection faces deployment challenges including data scarcity, multimodal alignment, reasoning reliability, and real-time constraints [51, 53].
 VII-A Data Scarcity
-LLM and VLM models require large, diverse datasets of labeled video-caption pairs to achieve robust performance in crash detection.
-However, such datasets are scarce, particularly for crash-specific scenarios, due to the rarity of crash events and the complexity of annotating video data [38].
-While recent large-scale datasets like SHRP 2 NDS [42] have significantly improved data availability, they still represent a fraction of the data needed to cover all possible real-world scenarios.
-Existing datasets, such as the Dashcam Accident Dataset (DAD) and UCF-Crime, contain limited crash-related videos, often with coarse or incomplete annotations [38, 41].
-For instance, DAD includes only 1,500 annotated videos, insufficient for training large-scale multimodal models like Video-LLaMA [27].
-Synthetic datasets, such as those used in CrashLLM, attempt to address this by generating video-caption pairs, but they may lack the variability of real-world traffic conditions [33].
-Data scarcity also exacerbates dataset imbalance, where non-crash videos dominate, leading to biased model predictions [11].
-Efforts to create comprehensive, crash-specific datasets are critical to improving model generalization and performance [39].
+Current crash datasets (DAD: 1,500 videos) are insufficient for large-scale multimodal models due to crash rarity and annotation complexity [61, 38].
+Solutions include synthetic data generation via CARLA/SUMO simulators, self-supervised pre-training on unlabeled driving data, and federated learning for privacy-preserving collaboration.
 VII-B Multimodal Alignment
-Achieving temporal synchronization between video frames and language descriptions is a significant challenge, particularly in crash detection scenarios involving occlusions, low framerates, or complex traffic dynamics [29].
-Multimodal alignment requires models to correlate visual events (e.g., a collision) with corresponding textual descriptions (e.g., “A car collided with a truck”), ensuring temporal and semantic consistency [27].
-For example, frameworks like VERA struggle to align frame-level visual features with textual captions under occlusions, where key crash indicators may be obscured [29].
-Low framerates, common in traffic surveillance videos, further complicate alignment by reducing temporal resolution, as noted in studies of classical optical flow methods [4].
-Cross-attention mechanisms, as used in VAD-LLaMA, mitigate this by modeling long-range dependencies, but they increase computational complexity [34].
-Developing efficient alignment techniques that handle occlusions and variable framerates remains an open research problem [11].
+Temporal synchronization between video and language is challenging under occlusions and variable framerates [54]. Cross-attention mechanisms address this but increase computational complexity [55].
 VII-C Reasoning and Explainability
-While LLMs generate fluent crash reports, their outputs may be factually incorrect or inconsistent with visual inputs, a phenomenon known as hallucination [30].
-This poses a critical challenge for crash detection, where accurate and interpretable outputs are essential for applications like autonomous driving and post-event analysis [33].
-For instance, CrashLLM’s generative descriptions may inaccurately attribute crash causes (e.g., “sudden braking” instead of “lane change”) if visual cues are ambiguous [33].
-Ensuring consistency between LLM-generated reports and visual evidence requires robust grounding techniques, such as those explored in VERA, which uses LLaVA-1.5 to verbalize crash rationales [29].
-However, even advanced models struggle with causal reasoning under complex traffic scenarios, as noted in studies of multimodal forecasting [35].
-Improving explainability through techniques like scene-graph grounding or causal intervention modules is a key open problem to enhance trust in LLM-based systems [26, 11].
+LLM hallucination risks factually incorrect crash reports in safety-critical applications [17].
+Robust grounding techniques and causal intervention modules are needed to ensure visual-textual consistency [54, 51].
 VII-D Real-time Constraints
-Real-time crash detection is critical for applications like accident response and autonomous vehicle decision-making, but LLM-based pipelines often exceed the computational limits required for real-time inference [37].
-Large models, such as GPT-4 or Video-LLaMA, demand significant resources, with inference times ranging from hundreds of milliseconds to seconds per video clip, far exceeding the 30–100 ms latency required for real-time systems [30, 27].
-For example, while the specific hardware is not detailed, the VADSK paper notes that the foundational models used within Holmes-VAD have inference speeds ranging from approximately 2 to 6 seconds per frame, clearly outside the real-time window for immediate response [58].
-This highlights a significant performance gap that needs to be addressed for practical deployment.
-Optimizing model architectures (e.g., using lightweight adapters like LoRA [32]), developing efficient temporal modules, and leveraging quantization techniques are essential to meet real-time constraints while maintaining performance [11].
-For example, Holmes-VAD’s multi-modal LLM achieves high accuracy but incurs high computational costs, limiting its use in resource-constrained environments [56].
-Training-free approaches like LAVAD reduce latency by eliminating fine-tuning, achieving 85.0% AUC with lower overhead, but they may sacrifice precision in complex scenarios [37].
-Classical methods, such as frame differencing, offer lower latency but lack the robustness of LLMs [3].
-A quantitative comparison of inference times and hardware requirements across models remains a challenge due to a lack of standardized reporting in the literature, but the available data suggests a clear trade-off between model complexity, accuracy, and processing speed.
+LLM-based models for video analysis often face inference latencies in the range of hundreds of milliseconds to seconds, which can exceed the 30–100 ms per-frame budget typically required for real-time safety-critical applications [73, 88].
+While exact timings are rarely reported in crash-specific studies, the computational overhead of 7B+ parameter models is widely acknowledged.
+Potential solutions discussed in the broader literature include model quantization for reduced compute cost, hierarchical pipelines where lightweight detectors filter candidate events before LLM-based reasoning, and edge–cloud hybrid architectures for distributing workloads [17, 87].
 VII-E Ethical and Fairness Considerations
-Beyond technical hurdles, the deployment of LLM-based crash detection systems raises significant ethical and fairness concerns.
-Datasets such as DAD and UCF-Crime, while foundational, may contain inherent biases by over-representing specific geographic regions, weather conditions, or vehicle types.
-Models trained on such data may perform poorly in underrepresented scenarios, leading to inequitable safety outcomes.
-Furthermore, the use of dashcam and surveillance footage introduces profound privacy risks, as this data can capture identifiable information like faces and license plates.
-The societal impact of large-scale video data collection and analysis requires careful consideration of data anonymization techniques and transparent policies regarding data usage and storage.
-Algorithmic fairness is another critical issue.
-Biased data can lead to biased models, where false positive or negative rates are disproportionately higher for certain demographic groups or environments.
-While methods like Holmes-VAD aim to address unbiased detection, the risk remains that automated systems could perpetuate or even amplify existing societal biases, with serious consequences for insurance liability and legal accountability.
-Therefore, developing fair, transparent, and privacy-preserving systems is not just a technical challenge but an ethical imperative for the responsible advancement of this technology.
-VII-F Robustness and Generalization Analysis
-A critical challenge for LLM-based crash detection models is their ability to generalize to scenarios not well represented in the training data.
-While models may perform well on benchmarks, their robustness in unseen or out-of-distribution (OOD) scenarios—such as rural roads, adverse weather (e.g., snow, heavy rain), or different geographic regions—remains a significant concern.
-This limitation poses a risk of overfitting to the visual and contextual cues prevalent in the training data, leading to a sharp degradation in performance when deployed in the real world [59].
-VII-F1 Domain Generalization
-To address this, domain generalization techniques are essential.
-These methods aim to learn domain-invariant features that are robust to shifts in data distribution.
-Techniques such as domain-invariant feature learning, meta-learning, and adversarial domain adaptation can help models generalize from multiple source domains to unseen target domains [60].
-Data augmentation, particularly using high-fidelity simulators like CARLA, allows for the creation of diverse training data by simulating varied weather and lighting conditions, which can significantly improve model robustness [61].
-VII-F2 Out-of-Distribution (OOD) Detection
-Complementary to domain generalization is Out-of-Distribution (OOD) detection, which focuses on identifying whether a given input is from a novel, unseen distribution.
-In crash detection, this is crucial for flagging rare events (e.g., unusual multi-vehicle pile-ups, crashes involving non-standard vehicles) that the model was not trained to recognize.
-By equipping models like LAVAD or VERA with an OOD detection mechanism, the system could defer to a human operator or a fail-safe mode when it encounters a scenario it cannot confidently analyze, thereby increasing safety.
-Methods based on uncertainty estimation are particularly promising for this task, as they can quantify the model’s confidence in its predictions [62].
-VII-F3 Cross-Dataset Evaluation and Quantitative Analysis
-A rigorous evaluation methodology is necessary to properly assess generalization.
-This should involve comprehensive cross-dataset evaluation, where models trained on one dataset (e.g., DAD) are tested on another with different characteristics (e.g., UCF-Crime or the more diverse SHRP 2 NDS) [40, 42].
-A quantitative analysis of this performance drop, for example, by reporting the decrease in AUC or F1-score, provides a clear measure of a model’s generalization capability.
-Future work should standardize this practice and ideally present results in a comparative table, analyzing performance drops with statistical significance tests to build a clearer picture of which models are most robust.
-VII-F4 Augmentation with Synthetic Data
+Dataset biases and privacy concerns require stratified data construction, bias auditing, differential privacy, and clear accountability frameworks with explainable AI for safety-critical applications.
+VII-F Robustness and Failure Analysis
+Models show significant performance drops (12-16%) in out-of-distribution scenarios [89].
+Critical Failure Modes: (1) Occlusion failures where 40% of crashes are missed when key vehicles are partially obscured, (2) Weather degradation with 25% accuracy drop in rain/snow conditions, (3) Lighting failures in dawn/dusk scenarios reducing detection by 30%, (4) Adversarial vulnerability where imperceptible perturbations can cause 60% false negatives, raising security concerns for malicious attacks.
+Solutions include domain generalization, adversarial training, OOD detection mechanisms, and systematic cross-dataset evaluation [90, 91].
+VII-F1 Augmentation with Synthetic Data
 The scarcity of diverse, real-world crash data is a major bottleneck.
@@
 By training or fine-tuning models on a combination of real and synthetic data, it is possible to significantly improve their ability to generalize to new environments and situations, making them more reliable for real-world deployment.
-The challenges of data scarcity, multimodal alignment, reasoning and explainability, real-time constraints, and generalization underscore the complexities of deploying LLM and VLM-based crash detection systems.
-Data scarcity limits model generalization, necessitating larger, more diverse datasets [39].
-Multimodal alignment issues highlight the need for advanced fusion strategies, such as cross-attention, to handle occlusions and low framerates [34].
-Reasoning and explainability challenges demand robust grounding techniques to ensure factual accuracy [29].
-Real-time constraints require optimized architectures to balance latency and performance [37].
-Finally, ensuring robust generalization requires a combination of advanced techniques like domain generalization, OOD detection, and the use of synthetic data, all validated through rigorous cross-dataset evaluation.
-Addressing these challenges will enhance the reliability and applicability of LLM-based crash detection, building on the advancements discussed in Section VI and paving the way for future research in intelligent transportation systems.
-VIII Ethical, Legal, and Societal Implications
-Beyond the technical challenges, the deployment of LLM-based crash detection systems carries profound ethical, legal, and societal implications that demand careful consideration.
-As these technologies become more integrated into safety-critical domains, a proactive approach to addressing their broader impact is essential for responsible innovation and public trust.
-VIII-A Bias and Fairness
-A primary concern is the risk of algorithmic bias.
-Datasets used to train crash detection models, such as DAD and UCF-Crime, often over-represent urban environments and specific weather conditions, potentially leading to biased models that underperform in rural or atypical scenarios [63].
-This can result in inequitable safety outcomes, where the system is less reliable for certain communities.
-To mitigate this, researchers propose debiasing techniques like reweighting underrepresented data samples and adversarial training to create more robust and equitable models.
-Ensuring fairness requires a concerted effort to build diverse and representative datasets that reflect the full spectrum of real-world driving conditions.
-VIII-B Privacy Risks
-The use of dashcam and surveillance footage for crash detection raises significant privacy concerns, as this data can capture personally identifiable information such as faces and license plates.
-The large-scale collection and analysis of such video data require robust privacy-preserving measures.
-Techniques like homomorphic encryption, which allows computation on encrypted data, and differential privacy, which adds statistical noise to data to protect individual identities, are promising solutions [64].
-Furthermore, strict adherence to data protection regulations like GDPR and CCPA is necessary to ensure that data is handled responsibly and transparently.
-VIII-C Legal Accountability
-The increasing autonomy of crash detection systems introduces complex questions of legal accountability.
-When an LLM-based system like CrashLLM or VERA generates a crash report that is used in insurance claims or legal proceedings, determining liability in the event of an error becomes challenging.
-Is the responsibility with the vehicle owner, the software developer, or the vehicle manufacturer?
-A clear legal framework is needed to address these new liability scenarios.
-Recent surveys on AI safety assurance emphasize the need for new standards and regulations that cover the entire lifecycle of AI systems, from design and testing to deployment and maintenance, to establish clear lines of accountability [65].
-VIII-D Societal Impact and Public Trust
-The societal impact of LLM-based crash detection includes significant benefits, such as the potential for faster emergency response and a reduction in road fatalities.
-However, there are also risks, including over-reliance on automated systems and public apprehension.
-Building public trust is paramount for the successful adoption of these technologies.
-Surveys on public trust in AI reveal that while citizens see the potential benefits, they also harbor significant concerns about job loss, privacy, and misuse of the technology by authorities [66].
-Fostering trust requires transparent communication about how the technology works, its limitations, and the safeguards in place.
-Engaging the public in the development and oversight process and ensuring that the systems are explainable and accountable will be key to building public confidence [63].
-VIII-E Case Study: A Hypothetical Scenario
-Consider a hypothetical scenario where an LLM-based crash detection system is widely deployed in a country.
-The system works exceptionally well in urban areas but has not been sufficiently trained on data from rural regions with unpaved roads and different types of agricultural vehicles.
-An accident occurs on a rural road involving a tractor, but the system fails to detect it in a timely manner, delaying emergency services.
-The legal investigation reveals the system’s biased performance.
-This scenario highlights the severe consequences of dataset bias, leading to legal challenges against the manufacturer for negligence, regulatory fines for deploying an unsafe product, and a significant erosion of public trust in AI-powered safety systems.
-It underscores the critical need for comprehensive testing, fair representation in training data, and robust regulatory oversight before widespread deployment.
-IX Integration with Autonomous Driving Ecosystems
+VII-G Domain Adaptation and Generalization Strategies
+Current models show significant performance drops (12-16% on average) when deployed in new environments. We propose several strategies to address this critical limitation:
+VII-G1 Geographic Adaptation
+Multi-region training includes data from diverse geographic regions (urban/rural, different countries) in training sets.
+Style transfer uses domain adaptation techniques to transfer models trained on one region to another.
+Local fine-tuning develops protocols for rapid adaptation using 100-500 local examples.
+VII-G2 Weather and Lighting Robustness
+Adversarial weather training augments training data with synthetic weather conditions (rain, snow, fog, night).
+Multi-spectral integration combines visible and infrared cameras for all-weather operation.
+Temporal consistency uses video sequences rather than single frames to maintain performance in poor visibility.
+VII-G3 Camera and Hardware Adaptation
+Resolution invariance trains models to handle varying camera resolutions and viewing angles.
+Calibration-free methods develop approaches that don’t require precise camera calibration.
+Hardware-agnostic deployment ensures models work across different edge computing platforms.
+These challenges reveal a significant deployment gap between laboratory results and real-world requirements, highlighting the fundamental mismatch between data-hungry LLMs and resource-constrained transportation systems [62, 60].
+VIII Integration with Autonomous Driving Ecosystems
+Figure 8: Integration of an LLM-based crash detection module within a standard Autonomous Vehicle (AV) perception-planning-action pipeline.
+The model provides semantic risk assessment to the planning module, enabling more nuanced and proactive vehicle control.
 The practical value of LLM-based crash detection is maximized when it is seamlessly integrated into the broader autonomous vehicle (AV) ecosystem.
@@
 This section explores how these models interface with core AV modules.
-IX-A AV Pipeline Integration
+VIII-A AV Pipeline Integration
 A standard AV software architecture follows a perception-planning-action pipeline [93].
 LLM-based crash detection serves as an advanced component of the perception stack.
-The outputs from a model like VERA or CrashLLM—such as a textual description (”imminent rear-end collision due to sudden braking”) or a classification of risk—can be fed directly into the AV’s decision-making or planning module.
+The outputs from a model like VERA or CrashLLM—such as a textual description (imminent rear-end collision due to sudden braking”) or a classification of risk—can be fed directly into the AV’s decision-making or planning module.
 This allows the AV to move beyond simple object detection and react to a nuanced understanding of the scene’s dynamics.
 For example, a high-risk prediction could trigger an emergency braking maneuver or a swerving action, demonstrating a direct link between semantic crash understanding and vehicle control.
-IX-B Multi-Sensor Fusion
+VIII-B Multi-Sensor Fusion
 While cameras are the primary input for the VLM-based methods discussed, their fusion with other sensor modalities like LiDAR and radar is crucial for building a robust perception system for AVs.
@@
 A multi-sensor fusion approach would combine the semantic, contextual understanding from an LLM/VLM with the geometric and velocity data from LiDAR/radar [94].
-For example, a VLM might identify a ”pedestrian looking at their phone,” and this semantic flag could be fused with LiDAR data confirming their trajectory towards the road, leading to a more reliable and timely pre-emptive braking action than either sensor could achieve alone.
+For example, a VLM might identify a “pedestrian looking at their phone,” and this semantic flag could be fused with LiDAR data confirming their trajectory towards the road, leading to a more reliable and timely pre-emptive braking action than either sensor could achieve alone.
 Extending frameworks like Video-LLaMA to incorporate these additional data streams is a promising research direction.
-IX-C Real-time Performance and Latency
+VIII-C Real-time Performance and Latency
 A critical requirement for any system integrated into an AV pipeline is real-time performance, typically with a latency of under 100 ms, and often as low as 30 ms for critical safety functions.
@@
 Lightweight models, such as MiniGPT-4 [68], are specifically designed for resource-constrained environments and offer a viable path for deploying advanced reasoning capabilities on embedded AV hardware.
-IX-D Crash Anticipation as a Proactive Safety Measure
+VIII-D Crash Anticipation as a Proactive Safety Measure
 Modern AV safety is shifting from reactive collision avoidance to proactive crash anticipation.
@@
 This proactive stance, informed by the rich contextual understanding of video data, represents a significant step towards achieving Level 4 and Level 5 autonomy.
-X Quantitative Benchmarking and Ablation Studies
-To provide a clearer, more rigorous comparison of the methods discussed, this section proposes a framework for unified quantitative benchmarking and outlines the need for systematic ablation studies.
-X-A Cross-Dataset Generalization
-A crucial test of robustness is cross-dataset generalization.
-Models should be trained on one dataset (e.g., CADP) and evaluated on another (e.g., BDD100K) without fine-tuning.
-This tests the model’s ability to generalize to new environments, camera types, and traffic patterns.
-Reporting the performance drop in such scenarios would provide a quantitative measure of a model’s generalization capabilities, which is currently a major gap in the literature.
-X-B Ablation Studies
-To understand the contribution of individual architectural components, systematic ablation studies are needed.
-For instance, in a model like VERA, one could ablate the language-based reasoning module to quantify its impact on performance compared to using only the visual features.
-Similarly, for Video-LLaMA, ablating the temporal module would reveal how much the model relies on temporal cues versus single-frame analysis.
-Such studies are essential for understanding what makes these models effective and for guiding future research [70].
-X-C Computational Metrics
-For practical deployment, especially in AVs, computational performance is as important as accuracy.
-Future benchmarking should include a systematic comparison of inference latency (in milliseconds) and computational requirements (e.g., GPU memory, TFLOPs) on standardized hardware.
-For example, comparing the performance of a heavyweight model like Holmes-VAD against a lightweight, training-free model like LAVAD would highlight the trade-offs between accuracy and efficiency, providing critical insights for real-world applications where resources are constrained [58].
-XI Future Directions
-The challenges outlined in Section VII and the advancements in LLM-based crash detection methods discussed in Section VI highlight the need for innovative approaches to enhance the robustness, scalability, and applicability of these systems in intelligent transportation systems.
-This section proposes key future directions to address current limitations and advance the field.
-XI-A Synthetic Training Data
-The scarcity of labeled video-caption pairs for crash detection, as noted in Section VII, limits the generalization of LLMs and VLMs [38, 39].
-To address this, synthetic training data generated through simulation platforms like CARLA or SUMO offers a promising solution [40].
-These platforms can create diverse crash scenarios, including rare events (e.g., multi-vehicle collisions under adverse weather), paired with detailed textual narratives (e.g., “A truck skidded on a wet road, colliding with a sedan”).
-For instance, CARLA’s high-fidelity simulations have been used to generate video data for autonomous driving tasks, which can be extended to crash-specific datasets with annotations for temporal events and causal factors [40].
-Synthetic data can augment existing datasets like DAD and UCF-Crime, reducing dataset imbalance and improving model robustness [41].
-Recent work on CrashLLM demonstrates the potential of synthetic video-caption pairs, though real-world variability remains a challenge [33].
-Future efforts should focus on hybrid datasets combining synthetic and real-world data, leveraging techniques like domain adaptation to ensure transferability [23].
-Such approaches can mitigate data scarcity, enabling models to handle diverse traffic conditions and rare crash scenarios [11].
-XI-B Video-grounded QA Benchmarks
+IX Future Directions
+Based on the challenges and advancements discussed, this section proposes key future directions to enhance robustness, scalability, and applicability of LLM-based crash detection systems.
+IX-A Synthetic Training Data
+Data scarcity limits LLM generalization [61, 62].
+Synthetic training data from CARLA/SUMO simulators can generate diverse crash scenarios with textual narratives, augmenting datasets like DAD and UCF-Crime [63, 64].
+Hybrid synthetic-real datasets with domain adaptation techniques offer scalable solutions for rare crash scenarios [53, 51].
+IX-B Video-grounded QA Benchmarks
 Current crash detection systems, such as VERA and Holmes-VAD, excel in generating descriptive crash reports but lack robust question-answering (QA) capabilities for interactive analysis [54, 85].
@@
 Creating such benchmarks will improve model interpretability and enable interactive crash analysis, addressing the reasoning challenges noted in Section VII [17].
-XI-C Fine-tuned VLMs
+IX-C Fine-tuned VLMs
 While pre-trained VLMs like BLIP-2 and Flamingo offer strong generalization, their performance in crash-specific scenarios is limited by the lack of domain-specific fine-tuning [19, 18].
@@
 Recent work on VERA demonstrates the benefits of fine-tuning LLaVA-1.5 for explainable crash detection, achieving high AUC and BLEU scores [54].
-Future efforts should focus on curating crash-specific annotations, incorporating temporal models like TimeSformer or SlowFast networks to capture dynamic events [52, 46].
-Such fine-tuned VLMs will enhance accuracy and robustness, particularly in challenging scenarios involving occlusions or low framerates, as discussed in Section VII [11].
-XI-D Multilingual and Low-resource Models
-Most LLM and VLM-based crash detection systems rely on English-language datasets and descriptions, limiting their applicability in non-English-speaking regions [33].
-Expanding to multilingual and low-resource models can improve accessibility, enabling crash detection systems to process descriptions in languages like Spanish, Mandarin, or Hindi.
-Multilingual pre-training, as explored in models like MiniGPT-4, provides a foundation for handling diverse linguistic inputs [45].
-For example, fine-tuning Video-LLaMA on multilingual crash datasets, such as those derived from global traffic surveillance, can enable cross-linguistic crash analysis [27].
-Low-resource models, optimized for environments with limited computational resources, are equally critical for deployment in developing regions.
-Techniques like knowledge distillation or training-free approaches, as used in LAVAD, can reduce model complexity while maintaining performance [37].
-Initiatives like TrafficLens demonstrate the potential of lightweight models for video-to-text analysis in resource-constrained settings [36].
-Future work should prioritize multilingual dataset creation and efficient model architectures, ensuring equitable access to crash detection technologies across diverse linguistic and economic contexts [11, 10].
-XI-E Edge Computing and Distributed Learning
-The real-time and privacy-sensitive nature of crash detection necessitates a shift towards edge computing and distributed learning architectures.
-These approaches address the latency and data privacy challenges inherent in cloud-centric models.
-XI-E1 Edge-Optimized Models
-Deploying complex VLMs on resource-constrained edge devices (e.g., in-vehicle computers) requires significant model optimization.
-Techniques like pruning (removing redundant model weights), quantization (reducing the numerical precision of weights), and knowledge distillation (training a smaller model to mimic a larger one) are crucial for reducing model size and latency [69].
-Research in edge-based video analytics for ITS has shown that such optimizations can enable real-time performance on embedded hardware [71].
-For example, a quantized and pruned version of a model like MiniGPT-4 could be deployed on edge cameras for real-time crash detection, as its smaller footprint would meet the strict computational budget of in-vehicle systems.
-XI-E2 Federated Learning for Privacy-Preserving Training
-Federated learning offers a powerful paradigm for training models on decentralized data sources, such as data from a fleet of vehicles or a network of traffic cameras, without compromising privacy.
-In this approach, the model is trained locally on each device, and only the model updates (gradients), not the raw data, are sent to a central server for aggregation [72].
-This is particularly valuable for crash detection, as it allows a global model to learn from a diverse range of driving scenarios and environments without requiring sensitive video data to be uploaded to the cloud, thus preserving the privacy of drivers.
-XI-E3 Distributed Inference for Low Latency
-For very large models, even after optimization, inference may be too slow for a single edge device.
-Distributed inference, where the computational load is split between the edge and the cloud, offers a solution.
-For example, an initial, low-latency analysis could be performed on the edge device to detect potential anomalies, and if a high-risk event is flagged, more detailed video data could be sent to a powerful cloud-based model for a more thorough analysis and descriptive summary [73].
-This hybrid approach balances the need for low latency with the desire for deep, semantic understanding of crash events.
-A case study could involve deploying a lightweight variant of LAVAD on edge cameras to continuously monitor for anomalies, with its outputs triggering a more powerful VLM in the cloud for detailed causal analysis, with latency being a key evaluation metric on a dataset like CCD [37, 44].
-XI-F Active Learning and Human-in-the-Loop (HITL)
-To further address data scarcity and improve model reliability, active learning and Human-in-the-Loop (HITL) systems present a promising frontier.
-These approaches can create a virtuous cycle of data collection, annotation, and model refinement, making the entire system more efficient and trustworthy.
-XI-F1 Active Learning for Efficient Annotation
-Active learning can significantly reduce the manual annotation burden by intelligently selecting the most informative data for labeling [74].
-For crash detection, an active learning system could prioritize video clips where the model is most uncertain about its predictions [75].
-For example, after an initial model is trained on a small subset of DAD or CADP, it could be used to score a large, unlabeled pool of traffic videos.
-The videos with the highest uncertainty would then be sent to human annotators.
-This targeted annotation strategy ensures that labeling efforts are focused on the most challenging and valuable examples, accelerating model improvement and reducing the cost of creating large-scale datasets like DAD or CADP [38, 39].
-XI-F2 Human-in-the-Loop for Validation and Refinement
-HITL systems are essential for safety-critical applications, where the cost of an error is high.
-In the context of crash detection, a HITL system would involve human experts validating the outputs of models like VERA or CrashLLM [76].
-This is particularly important for correcting ”hallucinations” or factual inaccuracies in LLM-generated explanations of crash events.
-By providing corrective feedback, human experts can iteratively refine the model’s performance and ensure its outputs are reliable and trustworthy.
-This process also generates a valuable dataset of human-verified crash descriptions, which can be used for further fine-tuning.
-Evaluation metrics for such a system could go beyond standard captioning scores like BLEU to include human agreement scores and measures of annotation efficiency, providing a more holistic assessment of the system’s performance.
-The proposed future directions—synthetic training data, video-grounded QA benchmarks, fine-tuned VLMs, and multilingual and low-resource models—address the core challenges of LLM-based crash detection systems.
-Synthetic data generation can mitigate data scarcity, building on datasets like BDD100K and CADP [40, 39].
-QA benchmarks will enhance explainability, complementing methods like VERA [29].
-Fine-tuned VLMs will improve multimodal alignment, leveraging architectures like TimeSformer [52].
-Multilingual and low-resource models will ensure global applicability, aligning with real-time needs [37].
-These directions, supported by foundational work in vision-language pre-training [23, 24, 28] and temporal modeling [51, 49], pave the way for robust, interpretable, and accessible crash detection systems, advancing the field beyond the methods discussed in Section VI [11].
-XII Conclusion
-The advent of Large Language Models (LLMs) and Vision-Language Models (VLMs) has transformed crash detection from video data, enabling high-level reasoning that bridges visual observations with linguistic descriptions.
-This survey comprehensively reviewed advancements in LLM-based crash detection from 2023 to 2024, covering models, datasets, taxonomies, challenges, and future directions, with the goal of guiding researchers toward building reliable and explainable multimodal systems for safety-critical applications in intelligent transportation systems [11, 33].
-By synthesizing recent developments, this work underscores the potential of LLMs to revolutionize traffic safety while identifying critical areas for further innovation.
-Section VI analyzed state-of-the-art methods, such as CrashLLM, VERA, Holmes-VAD, and Video-LLaMA, which leverage multimodal fusion strategies (e.g., early, late, cross-attention) to achieve high accuracy and explainability [33, 29, 56, 27].
-These methods outperformed classical approaches, such as CNN-based systems, by integrating visual and textual data to address complex scenarios involving occlusions and variable lighting [3, 6].
-For instance, CrashLLM’s generative capabilities produce detailed crash descriptions, achieving over 92% accuracy, while VERA’s verbalized learning enhances post-event analysis with 86.55% AUC [33, 29].
-The survey also introduced a taxonomy of fusion strategies and LLM roles (e.g., passive, generative), providing a structured framework for understanding these advancements [11, 27].
-Key datasets, including BDD100K, CADP, and UCF-Crime, have been instrumental in training these models, though their limitations—such as limited crash-specific videos and coarse annotations—pose challenges [40, 39, 41].
-The recent emergence of large-scale, finely annotated datasets like SHRP 2 NDS [42] is a critical step forward.
-Section VII detailed these obstacles, highlighting data scarcity, multimodal alignment difficulties, reasoning errors (e.g., hallucination), and real-time computational constraints [38, 30, 37].
-For example, the Dashcam Accident Dataset (DAD) contains only 1,500 videos, insufficient for large-scale VLM training, while real-time systems like Video-LLaMA struggle with latency requirements of 30–100 ms [38, 27].
-These challenges underscore the need for innovative solutions to enhance robustness and practicality.
-Section XI proposed four directions to address these issues: generating synthetic training data using simulators like CARLA, developing video-grounded QA benchmarks, fine-tuning VLMs for crash-specific scenarios, and creating multilingual and low-resource models [40, 32, 45].
-These strategies aim to mitigate data scarcity, improve temporal alignment, enhance explainability, and ensure global accessibility.
-For instance, synthetic data can augment datasets like CADP, while QA benchmarks can enable interactive crash analysis [39].
-Fine-tuning with techniques like LoRA and multilingual pre-training with models like MiniGPT-4 can further advance performance and equity [32, 45].
-The implications of these advancements are profound for safety-critical applications, such as autonomous driving, traffic surveillance, and accident response.
-Methods like Holmes-VAD and LAVAD demonstrate the potential for unbiased and training-free crash detection, while TrafficVLM highlights the power of dense captioning for post-event analysis [56, 37, 43].
-Foundational models, including CLIP, Flamingo, and GPT-4, have laid the groundwork for robust feature extraction and reasoning, supported by evaluation metrics like BLEU, METEOR, and CIDEr [23, 26, 30, 53, 54, 55].
-Temporal modeling architectures, such as TimeSformer, SlowFast, and ConvLSTM, further enhance video understanding, addressing dynamic crash scenarios [52, 46, 51].
-This survey encourages researchers to build on these foundations, leveraging insights from recent methods and proposed directions to develop multimodal systems that are accurate, interpretable, and deployable in real-world settings.
-By addressing challenges through innovations like synthetic data and fine-tuned VLMs, the field can move toward standardized benchmarks and efficient architectures, as seen in works like CLIP4Clip and scaling approaches [28, 24].
-Ultimately, we hope this survey serves as a catalyst for advancing LLM-based crash detection, fostering safer and more intelligent transportation systems [11, 10].
+Future efforts should focus on curating crash-specific annotations, incorporating temporal models like TimeSformer or SlowFast networks to capture dynamic events.
+X Conclusion
+This survey examined LLM/VLM-based video crash detection and highlighted the persistent gap between promising research results and real-world deployment constraints.
+Recent systems show strong capabilities in multimodal understanding, but many works still underreport latency, struggle with cross-domain robustness, and rely on datasets that are relatively small or weakly annotated for safety-critical use.
+See, e.g., LLM/VLM methods such as Video-LLaMA for video understanding, and training-free approaches like LAVAD for anomaly detection [38, 60].
+X-A Key Findings
+  1. 1.
+    Deployment Gap: High-capacity VLMs can achieve strong accuracy on curated benchmarks, but real-time, edge-friendly latency is rarely reported and remains a major barrier for on-device ITS deployment.
+Training-free pipelines (e.g., LAVAD) reduce data/engineering overhead but depend on captioning backends whose speed varies across hardware and models [38, 60].
+  2. 2.
+    Generalization Challenge: Across the broader video anomaly literature, models often degrade when moved off their source domain/dataset; works frequently note domain specificity and the need for adaptation when transferring between datasets such as UCF-Crime, CADP, and DAD [10, 62, 61].
+  3. 3.
+    Data Bottlenecks: Several widely used datasets for crash/anomaly detection remain modest in size.
+For example, DAD contains $\sim$620 dashcam videos (with positive/negative splits), and UCF-Crime has $\sim$1,900 untrimmed surveillance videos across 13 anomaly classes [61, 10].
+X-B Critical Recommendations
+For Researchers: Design efficiency-first architectures and report latency alongside accuracy; incorporate domain adaptation/evaluation across datasets; and consider training-free or weakly-supervised regimes where appropriate.
+Privacy-preserving and federated paradigms are increasingly relevant for multi-site video data [96].
+For Practitioners: Start with lightweight or training-free baselines (e.g., LAVAD-style pipelines) before committing to heavy retraining; adopt hierarchical processing (quick filters $\to$ deeper reasoning on suspected segments); plan for local adaptation and monitoring; and keep human oversight in the loop for safety-critical decisions [60].
+X-C Synthesis
+We observe: (i) Reasoning vs.
+Deployability: richer temporal/contextual reasoning often coincides with higher compute cost; (ii) Source–Target Drift: models tuned to one dataset typically require adaptation to new camera placements, traffic patterns, and artifact distributions; (iii) System Design Tiers: in practice, teams converge on edge-first filters with selective offloading to stronger backends when latency budgets permit.
+These are qualitative patterns—reporting and standardized latency benchmarks remain an open need [38, 96].
+X-D The Path Forward
+Progress hinges on (a) stronger temporal modeling under efficiency constraints (e.g., video transformers like TimeSformer/ViViT), (b) robust multilingual and low-resource operation via VLM alignment/fine-tuning, (c) privacy-aware training via federated or distributed learning where data cannot leave premises, and (d) human-in-the-loop active learning for targeted data curation and assurance [38, 96].